### v1

# A rigidity theorem for complex Kleinian groups

Abstract.
Farre, Pozzetti and Viaggi proved that any (dk)(d-k)-hyperconvex subgroup of $\mathsf{PSL}(d,\mathbb{C})$ is virtually isomorphic to a convex cocompact Kleinian group and that its $\alpha_{k}$-critical exponent is at most 2. We show that a (dk)(d-k)-hyperconvex subgroup is isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$ if and only if its $\alpha_{k}$-critical exponent is exactly 2. Furthermore, we show that if a strongly irreducible (dk)(d-k)-hyperconvex subgroup has $\alpha_{k}$-critical exponent $2$, then it is the image of a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$ by an irreducible representation of $\mathsf{PSL}(2,\mathbb{C})$ into $\mathsf{PSL}(d,\mathbb{C})$.
Canary was partially supported by grant DMS-2304636 from the National Science Foundation. Zhang was partially supported by the NUS-MOE grants A-8001950-00-$0$ and A-8000458-00-$0$. Zimmer was partially supported by a Sloan research fellowship and grants DMS-2105580 and DMS-2452068 from the National Science Foundation.
Dedicated to Pepe Seade on the occasion of his seventieth birthday
1. Introduction
In this paper, we continue the study of (dk)(d-k)-hyperconvex subgroups of $\mathsf{PSL}(d,\mathbb{C})$ initiated by Pozzetti–Sambarino–Wienhard [30] and Farre–Pozzetti–Viaggi [19]. A subgroup of $\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex for some $k\in\{1,\ldots,d-1\}$ if it is Anosov with respect to the simple roots $\alpha_{k-1}$, $\alpha_{k}$ and $\alpha_{k+1}$ and triples of points in its limit set satisfy a specific transversality condition (see Section 2.5 for the definition). The most basic examples of $k$-hyperconvex groups are obtained by considering the image of a convex cocompact subgroup of $\mathsf{PSL}(2,\mathbb{C})$ by an irreducible representation of $\mathsf{PSL}(2,\mathbb{C})$ into $\mathsf{PSL}(d,\mathbb{C})$.
Farre–Pozzetti–Viaggi showed that any (dk)(d-k)-hyperconvex subgroup of $\mathsf{PSL}(d,\mathbb{C})$ is virtually isomorphic to a convex cocompact subgroup of $\mathsf{PSL}(2,\mathbb{C})$. They further showed that the $\alpha_{k}$-critical exponent $\delta^{\alpha_{k}}(\Gamma)$ of a $k$-hyperconvex subgroup $\Gamma$ of $\mathsf{PSL}(d,\mathbb{C})$ is at most 2. Recall that the $\alpha_{k}$-critical exponent $\delta^{\alpha_{k}}(\Gamma)$ is the exponential growth rate of the number of elements $\gamma\in\Gamma$ so that
$\alpha_{k}(\kappa(\gamma))=\log\frac{\sigma_{k}(\gamma)}{\sigma_{k+1}(\gamma)}\leq T,$
where $\sigma_{1}(\gamma)\geq\cdots\geq\sigma_{d}(\gamma)$ are the singular values of $\gamma$.
It then follows from work of Pozzetti–Sambarino–Wienhard [30], that if $\Gamma$ is (dk)(d-k)-hyperconvex, then the Hausdorff dimension of the limit set $\Lambda_{k}(\Gamma)$ of the action of $\Gamma$ on the Grassmanian $\mathrm{Gr}_{k}(\mathbb{C}^{d})$ is exactly $\delta^{\alpha_{k}}(\Gamma)\leq 2$. If $\Gamma$ is virtually isomorphic to a lattice in $\mathsf{PSL}(2,\mathbb{C})$, then $\Lambda_{k}(\Gamma)$ is homeomorphic to a two-sphere, so we must have $\delta^{\alpha_{k}}(\Gamma)=2$. In this paper, we prove that the converse also holds.
Theorem 1.1.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, then $\delta^{\alpha_{k}}(\Gamma)=2$ if and only if $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$.
One may view this as a generalization of the fact, due to Sullivan [34] and Tukia [36], that the critical exponent of a convex cocompact subgroup of $\mathsf{PSL}(2,\mathbb{C})$ has Hausdorff dimension two if and only if it is a uniform lattice.
The only known examples where $\Gamma$ is (dk)(d-k)-hyperconvex, strongly irreducible and virtually isomorphic to a lattice in $\mathsf{PSL}(2,\mathbb{C})$ are obtained from lattices in $\mathsf{PSL}(2,\mathbb{C})$ by applying an irreducible representation of $\mathsf{PSL}(2,\mathbb{C})$ into $\mathsf{PSL}(d,\mathbb{C})$. We show that these are the only examples.
Theorem 1.2.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, strongly irreducible and $\delta^{\alpha_{k}}(\Gamma)=2$, then there exist a uniform lattice $\Gamma^{\prime}\subset\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ and an irreducible representation $\tau:\mathsf{PSL}(2,\mathbb{C})\to\mathsf{PSL}(d,\mathbb{C})$ so that $\Gamma=\tau(\Gamma^{\prime})$.
Since the property of being (dk)(d-k)-hyperconvex is stable (see [30, Prop. 6.2]), one may view our result as a global analogue of the local rigidity result of Raghunathan [31] (see also Menal–Ferrer and Porti [26, 29]) who proved that if $\Gamma$ is a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, then $\tau|_{\Gamma}$ is locally rigid in the character variety ${\bf X}(\Gamma,\mathsf{PSL}(d,\mathbb{C}))$ if $\tau:\mathsf{PSL}(2,\mathbb{C})\to\mathsf{PSL}(d,\mathbb{C})$ is an irreducible representation.
As a corollary to Theorem 1.2 we have the following.
Corollary 1.3.
If $\Gamma^{\prime}\subset\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is a uniform lattice and $\rho:\Gamma^{\prime}\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ is a strongly irreducible representation with finite kernel and (dk)(d-k)-hyperconvex image, then $\rho$ extends to an irreducible representation $\mathsf{PSL}(2,\operatorname{\mathbb{C}})\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{C}})$.
Any uniform lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is strongly irreducible and 1-hyperconvex, hence Corollary 1.3 can be viewed as a generalization of Mostow’s rigidity theorem [28] for hyperbolic 3-manifolds.
We also note that without irreducibility, there is considerable less rigidity. For instance, fix a uniform lattice $\Gamma_{0}\subset\mathsf{SL}(2,\operatorname{\mathbb{C}})$ and any representation $\rho:\Gamma_{0}\rightarrow\mathsf{SL}(d_{0},\operatorname{\mathbb{C}})$. Next let $\tau_{d}:\mathsf{SL}(2,\operatorname{\mathbb{C}})\rightarrow\mathsf{SL}(d,\operatorname{\mathbb{C}})$ denote the standard irreducible representation and define $\rho_{d}:\Gamma_{0}\rightarrow\mathsf{PSL}(d+d_{0},\operatorname{\mathbb{C}})$ by
$\rho_{d}(\gamma)=\begin{bmatrix}\rho(\gamma)&\\ &\tau_{d}(\gamma)\end{bmatrix}$.
Then for any $k\in\operatorname{\mathbb{N}}$ and $d$ sufficiently large (depending on $\rho$ and $k$) the image $\Gamma:=\rho_{d}(\Gamma_{0})$ is (dk)(d-k)-hyperconvex and has $\delta^{\alpha_{k}}(\Gamma)=2$.
The context
[…]
Another natural context is the theory of complex Kleinian groups which was developed by José Seade and his collaborators. A complex Kleinian group is a discrete subgroup of $\mathsf{PSL}(d,\mathbb{C})$. The theory of complex Kleinian groups developed independently and contemporaneously with the theory of Anosov representations. It seems to us that the stage is now set for these two fields to interact profitably. We recommend that book [13] as an introduction to the field of complex Kleinian groups.
Outline of proofs
The key tool in the proof of Theorem 1.1 is a deeper analysis of the limit sets of hyperconvex groups. We use the technology developed in [11] and [12] to show that if $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, then the $\alpha_{k}$-Patterson-Sullivan measure $\mu$ on $\Lambda_{k}(\Gamma)$ is $\delta:=\delta^{\alpha_{k}}(\Gamma)$-regular, i.e. there exist $D\geq 1$ and $t_{0}>0$ such that for every $x\in\Lambda_{k}(\Gamma)$ and every $0<t<t_{0}$, we have
$\frac{1}{D}t^{\delta}\leq\mu(B(x,t))\leq Dt^{\delta},$
where $B(x,t)$ is the open ball of radius $t$ about $x$ (see Theorem 3.1). Dey–Kim–Oh [18], see also Dey–Kapovich[17], previously showed that if $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov, $\phi\in\mathfrak{a}_{\theta}^{*}$ is symmetric and $\delta^{\phi}(\Gamma)<\infty$, then the $\phi$-Patterson-Sullivan measure on $\Lambda_{\theta}(\Gamma)$ is $\delta^{\phi}(\Gamma)$-Ahlfors regular with respect to a Finsler premetric on $\mathcal{F}_{\theta}$. In our case, $\alpha_{k}$ is not symmetric and the Ahlfors regularity is with respect to the Riemannian metric on $\mathrm{Gr}_{k}(\mathbb{C}^{d})$.
The following result is a simple consequence of Ahlfors regularity.
Theorem 1.4.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, then $\Lambda_{k}(\Gamma)$ has finite, non-zero $\delta^{\alpha_{k}}(\Gamma)$-dimensional Hausdorff measure.
Farre–Pozzetti–Viaggi [19, Thm. B] showed that if $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex and $\Gamma$ is not virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, then the limit set $\Lambda_{k}(\Gamma)$ of $\Gamma$ has 2-dimensional Hausdorff measure zero. Theorem 1.1 follows immediately from their result and Theorem 1.4.
We note that in the proof of Theorem 1.4 one may use standard properties of the exterior power representation to reduce the general case to the case where $k=1$.
One key input in the proof of Theorem 1.2 is the fact, which follows from work of Pozzetti–Sambarino–Wienhard [30] (see also Zhang–Zimmer [37]), that if $\Gamma$ is (dk)(d-k)-hyperconvex and $\partial\Gamma$ is a two-sphere, then $\Lambda_{k}(\Gamma)$ is a $\mathcal{C}^{1}$-smooth submanifold. We then observe that $\Lambda_{k}(\Gamma)$ is a complex submanifold of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$, and hence, by Chow’s theorem [14], a complex algebraic subvariety of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$. This implies that the real Zariski closure $\mathsf{G}$ of $\Gamma$ acts by biholomorphisms on $\Lambda_{k}(\Gamma)$. Then, since $\Lambda_{k}(\Gamma)$ is a two-sphere and hence biholomorphic to $\mathbb{CP}^{1}$, we obtain a surjective homomorphism from $\mathsf{G}$ to $\mathsf{PSL}(2,\mathbb{C})$, the biholomorphism group of $\mathbb{CP}^{1}$.
In Section 4 we establish a sufficient condition for the real Zariski closure of an Anosov group to be simple. Using this, we can prove that the homomorphism $\mathsf{G}\rightarrow\mathsf{PSL}(2,\mathbb{C})$ in the proof of Theorem 1.2 is indeed an isomorphism.
2. Background
2.1. The Jordan and Cartan projections
Let $\operatorname{\mathbb{K}}$ be either the field of real or complex numbers, and let $(e_{1},\dots,e_{d})$ be the standard basis of $\operatorname{\mathbb{K}}^{d}$. Recall that the standard Cartan subspace for $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, denoted $\mathfrak{a}$, is the real vector space of real-valued, $d\times d$, diagonal, traceless matrices:
$\mathfrak{a}=\{\mathrm{diag}(A_{1},\ldots,A_{d})\in\mathfrak{sl}(d,\operatorname{\mathbb{R}})\ |\ A_{1}+\cdots+A_{d}=0\}.$
For each $k\in\Delta:=\{1,\dots,d-1\}$, the $k$-th restricted simple root is the linear map
$\alpha_{k}:\mathfrak{a}\to\operatorname{\mathbb{R}}$
given by $\alpha_{k}(A)=A_{k}-A_{k+1}$. The standard positive Weyl chamber is the subset where all the simple roots are non-negative:
$\mathfrak{a}^{+}=\{A\in\mathfrak{a}\ |\ \alpha_{k}(A)\geq 0\text{ for all }k\in\Delta\}.$
Given $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, let
$\lambda_{1}(g)\geq\dots\geq\lambda_{d}(g)$
denote the modulus of the generalized eigenvalues of some lift $\tilde{g}$ of $g$ in $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$, and let
$\sigma_{1}(g)\geq\dots\geq\sigma_{d}(g)$
denote the singular values of $\tilde{g}$ (with respect to the standard Hermitian inner product on $\operatorname{\mathbb{K}}^{d}$). The Jordan projection and Cartan projection
$\nu,\kappa:\mathsf{PSL}(d,\mathbb{K})\to\mathfrak{a}^{+}$
are respectively given by
$\nu(g)=\mathrm{diag}(\log\lambda_{1}(g),\ldots,\log\lambda_{d}(g))\quad\text{and}\quad\kappa(g)=\mathrm{diag}(\log\sigma_{1}(g),\ldots,\log\sigma_{d}(g)).$
Recall that (see [4, Cor. 6.34])
$\lambda(g)=\lim\frac{1}{n}\kappa(g^{n}).$
By the singular value decomposition theorem, any $g\in\mathsf{PSL}(d,\mathbb{K})$ can be decomposed as $g=m_{1}\exp(\kappa(g))m_{2}$ for some $m_{1},m_{2}\in\mathsf{PSU}(d,\mathbb{K})$. Furthermore, if $k\in\{1,\dots,d-1\}$ satisfies $\alpha_{k}(\kappa(g))>0$, then
$U_{k}(g):=m_{1}(\operatorname{Span}_{\operatorname{\mathbb{K}}}(e_{1},\ldots,e_{k}))\in\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$
is well-defined (even though $m_{1}$ and $m_{2}$ are not uniquely specified by $g$). In this case, one can verify that $U_{d-k}(g^{-1})$ is also well-defined, and can be described as
$U_{d-k}(g^{-1})=m_{2}^{-1}(\operatorname{Span}_{\operatorname{\mathbb{K}}}(e_{k+1},\ldots,e_{d})).$
2.2. Flag manifolds and the Iwasawa decomposition
Given a non-empty subset
$\theta=\{k_{1},k_{2},\dots,k_{r}\}\subset\Delta,$
the $\theta$-flag manifold, denoted $\operatorname{\mathcal{F}}_{\theta}=\operatorname{\mathcal{F}}_{\theta}(\operatorname{\mathbb{K}}^{d})$, is the set of nested subspaces $x^{k_{1}}\subset\dots\subset x^{k_{r}}\subset\operatorname{\mathbb{K}}^{d}$ so that $x^{k_{j}}$ has dimension $k_{j}$ for each $j$. We often denote $x:=(x^{k_{1}},\dots,x^{k_{r}})$. Let $\operatorname{\mathsf{P}}_{\theta}\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ be the stabilizer of the flag $x_{0}\in\operatorname{\mathcal{F}}_{\theta}$ defined by $x_{0}^{k}=\operatorname{Span}(e_{1},\dots,e_{k})$ for all $k\in\theta$, where $e_{1},\dots,e_{d}$ is the standard basis of $\operatorname{\mathbb{K}}^{d}$. Since $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ acts transitively on the space of flags, we have the identification $\mathsf{PSL}(d,\operatorname{\mathbb{K}})/\mathsf{P}_{\theta}\cong\operatorname{\mathcal{F}}_{\theta}$ as $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$-spaces.
Let $\iota:\Delta\to\Delta$ be the map given by $\iota(k)=d-k$. We say that flags $x\in\operatorname{\mathcal{F}}_{\theta}$ and $y\in\operatorname{\mathcal{F}}_{\iota(\theta)}$ are transverse if $x^{k}+y^{d-k}=\operatorname{\mathbb{K}}^{d}$ for all $k\in\theta$. For every $x\in\operatorname{\mathcal{F}}_{\theta}$, let $\mathcal{Z}_{x}\subset\operatorname{\mathcal{F}}_{\iota(\theta)}$ denote the set of flags that are not transverse to $x$.
If $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ satisfies $\alpha_{k}(\kappa(g))>0$ for all $k\in\theta$, then we may define $U_{\theta}(g)\in\operatorname{\mathcal{F}}_{\theta}$ by $U_{\theta}(g)^{k}:=U_{k}(g)$ for all $k\in\theta$. The following is standard (see for example, [12, Prop. 2.3]).
Proposition 2.1.
Let $\theta\subset\Delta$ be non-empty. Suppose $x^{+}\in\operatorname{\mathcal{F}}_{\theta}$, $x^{-}\in\operatorname{\mathcal{F}}_{\iota(\theta)}$ and $(g_{n})_{n=1}^{\infty}$ is a sequence in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$. Then the following are equivalent:
  1. (1)
    $\alpha_{k}(\kappa(g_{n}))\to\infty$ for every $k\in\theta$, $U_{\theta}(g_{n})\rightarrow x^{+}$ and $U_{\iota(\theta)}(g_{n}^{-1})\rightarrow x^{-}$.
  2. (2)
    $g_{n}(x)\to x^{+}$ for all $x\in\operatorname{\mathcal{F}}_{\theta}\setminus\mathcal{Z}_{x^{-}}$, and this convergence is uniform on compact subsets of $\operatorname{\mathcal{F}}_{\theta}\setminus\mathcal{Z}_{x^{-}}$.
  3. (3)
    $g_{n}^{-1}(x)\to x^{-}$ for all $x\in\operatorname{\mathcal{F}}_{\iota(\theta)}\setminus\mathcal{Z}_{x^{+}}$, and this convergence is uniform on compact subsets of $\operatorname{\mathcal{F}}_{\iota(\theta)}\setminus\ \mathcal{Z}_{x^{+}}$.
  4. (4)
    There are open sets $\mathcal{U}^{+}\subset\operatorname{\mathcal{F}}_{\theta}$ and $\mathcal{U}^{-}\subset\operatorname{\mathcal{F}}_{\iota(\theta)}$ such that $g_{n}(x)\to x^{+}$ for all $x\in\mathcal{U}^{+}$ and $g_{n}^{-1}(x)\to x^{-}$ for all $x\in\mathcal{U}^{-}$.
By the Iwasawa decomposition theorem, we may decompose every $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ uniquely as $g=me^{A}n$ for some $m\in\mathsf{PSU}(d,\operatorname{\mathbb{K}})$, $A\in\mathfrak{a}$, and $n$ a unipotent upper triangular matrix in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$. Since $\mathsf{PSU}(d,\operatorname{\mathbb{K}})$ acts transitively on $\operatorname{\mathcal{F}}_{\Delta}$, we may use this to define the Iwasawa cocycle
$B:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\times\operatorname{\mathcal{F}}_{\Delta}\to\mathfrak{a}$
as follows. Given any $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ and any $x\in\operatorname{\mathcal{F}}_{\Delta}$, choose $k\in\mathsf{PSU}(d,\operatorname{\mathbb{K}})$ such that $x=k\operatorname{\mathsf{P}}_{\Delta}$. Then decompose $gk=me^{A}n$. It is straightforward to verify that $A$ does not depend on the choice of $k$, so we may define $B(g,x):=A$.
2.3. Divergent, transverse and Anosov subgroups
For any non-empty $\theta\subset\Delta$, a subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-divergent if for any $k\in\theta$ and any infinite sequence $\{\gamma_{i}\}$ in $\Gamma$, we have
$\lim_{i\to\infty}\alpha_{k}(\kappa(\gamma_{i}))=\infty.$
For any $\mathsf{P}_{\theta}$-divergent subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, we may define its $\theta$-limit set
$\Lambda_{\theta}(\Gamma):=\left\{\lim_{i\to\infty}U_{\theta}(\gamma_{i})\in\operatorname{\mathcal{F}}_{\theta}:(\gamma_{i})_{i=1}^{\infty}\text{ is an infinite sequence in }\Gamma\right\}.$
One can verify that $\Lambda_{\theta}(\Gamma)$ is a $\Gamma$-invariant subset of $\operatorname{\mathcal{F}}_{\theta}$, see [11, Prop. 3.1]. If $\theta$ is symmetric, i.e. $\theta=\iota(\theta)$, a $\mathsf{P}_{\theta}$-divergent subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is called $\mathsf{P}_{\theta}$-transverse if $\Lambda_{\theta}(\Gamma)\subset\operatorname{\mathcal{F}}_{\theta}$ is a transverse subset, i.e. any pair of distinct flags in $\Lambda_{\theta}(\Gamma)$ are transverse.
Recall that the action of a group $\Gamma$ by homeomorphisms on a compact metrizable topological space $M$ is called a convergence group action if for any infinite sequence $(\gamma_{i})_{i=1}^{\infty}$ in $\Gamma$, there is a subsequence $(\gamma_{i_{k}})_{k=1}^{\infty}$ and a pair of points $a,b\in M$ such that $\gamma_{i_{k}}(x)\to a$ for all $x\in M-\{b\}$. If $\Gamma$ acts on $M$ as a convergence group, a point $x\in M$ is a conical point if there exists a sequence $(\gamma_{i})_{i=1}^{\infty}$ in $\Gamma$ and a distinct pair of points $a,b\in M$ such that $\gamma_{i}(x)\to a$ and $\gamma_{i}(y)\to b$ for all $y\in M-\{x\}$.
Using Proposition 2.1, one can verify that if $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-transverse, then it acts on $\Lambda_{\theta}(\Gamma)$ as a convergence group, see [11, Prop. 3.3] or [23, Sec. 5.2]. We say that a $\mathsf{P}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-Anosov if every point in $\Lambda_{\theta}(\Gamma)$ is conical. Equivalently, a $\operatorname{\mathsf{P}}_{\theta}$-transverse subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\mathsf{P}_{\theta}$-Anosov if and only if $\Gamma$ is word hyperbolic and there exists a $\Gamma$-equivariant homeomorphism between $\Lambda_{\theta}(\Gamma)$ and the Gromov boundary $\partial\Gamma$ of $\Gamma$.
To avoid cumbersome notation, we often say a group is $\operatorname{\mathsf{P}}_{k}$-Anosov instead of $\operatorname{\mathsf{P}}_{\{k,d-k\}}$-Anosov.
The following is a characterization of Anosov representations due to Kapovich–Leeb–Porti [24, Thm. 1.5] and Bochi–Potrie–Sambarino [6, Prop. 4.6 and 4.9].
Theorem 2.2.
A subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is $\operatorname{\mathsf{P}}_{k}$-Anosov if and only if it is finitely generated and there exists $a,C>0$ such that
$\alpha_{k}(\kappa(\gamma))\geq a|\gamma|-C$
for all $\gamma\in\Gamma$, where $|\gamma|$ is the word length $\gamma$ with respect to some fixed finite generating set for $\Gamma$.
2.4. Patterson–Sullivan measures
For each $k=1,\dots,d-1$, the $k$-th restricted fundamental weight is the linear map
$\omega_{k}:\mathfrak{a}\to\operatorname{\mathbb{R}}$
given by $\omega_{k}(A)=A_{1}+\dots+A_{k}$. For any non-empty $\theta\subset\{1,\dots,d-1\}$, let
$\mathfrak{a}_{\theta}:=\{A\in\mathfrak{a}:\alpha_{k}(A)=0\text{ for all }k\notin\theta\}\quad\text{and}\quad\mathfrak{a}_{\theta}^{+}:=\mathfrak{a}^{+}\cap\mathfrak{a}_{\theta}.$
Observe that $\{\omega_{k}|_{\mathfrak{a}_{\theta}}:k\in\theta\}$ is a basis for $\mathfrak{a}_{\theta}^{*}$, so we may identify $\mathfrak{a}_{\theta}^{*}$ as the subspace of $\mathfrak{a}^{*}$ spanned by $\{\omega_{k}:k\in\theta\}$. We may then define a projection
$p_{\theta}:\mathfrak{a}\to\mathfrak{a}_{\theta}$
defined by $\omega_{k}(A)=\omega_{k}(p_{\theta}(A))$ for all $k\in\theta$. With this, define the $\theta$-partial Cartan projection to be the map
$\kappa_{\theta}:=p_{\theta}\circ\kappa:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\to\mathfrak{a}_{\theta}^{+}.$
Via the identification of $\mathfrak{a}_{\theta}^{*}$ as a subspace of $\mathfrak{a}^{*}$, for all $\phi\in\mathfrak{a}_{\theta}^{*}$ and $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, we have
$\phi(\kappa(g))=\phi(\kappa_{\theta}(g)).$
For all $k\in\Delta$, let $\left\|\cdot\right\|_{k}$ denote the norm on $\wedge^{k}\operatorname{\mathbb{K}}^{d}$ induced by the standard inner product on $\operatorname{\mathbb{K}}^{d}$. One may then verify that for all $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ and $x\in\operatorname{\mathcal{F}}_{\Delta}$, we have
$\omega_{k}(B(g,x))=\log\frac{\left\|g(v_{1})\wedge\dots\wedge g(v_{k})\right\|_{k}}{\left\|v_{1}\wedge\dots\wedge v_{k}\right\|_{k}},$
where $v_{1},\dots,v_{k}$ is some (any) basis of $x^{k}$. In particular, $\omega_{k}(B(g,x))$ depends only on $g$ and $x^{k}$. Thus, for any non-empty $\theta\subset\Delta$, we may define the $\theta$-partial Iwasawa cocycle
$B_{\theta}:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\times\operatorname{\mathcal{F}}_{\theta}\to\mathfrak{a}_{\theta}$
by $B_{\theta}(g,x)=p_{\theta}(B(g,\sigma(x)))$, where $\sigma:\operatorname{\mathcal{F}}_{\theta}\to\operatorname{\mathcal{F}}_{\Delta}$ is some (any) right inverse of the forgetful projection $\operatorname{\mathcal{F}}_{\Delta}\to\operatorname{\mathcal{F}}_{\theta}$. Again, via the identification of $\mathfrak{a}_{\theta}^{*}$ as a subspace of $\mathfrak{a}^{*}$, for all $\phi\in\mathfrak{a}_{\theta}^{*}$, $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, and $x\in\operatorname{\mathcal{F}}_{\theta}$, we have
$\phi(B_{\theta}(g,x))=\phi(B(g,\sigma(x)).$
Let $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ be $\mathsf{P}_{\theta}$-divergent and let $\phi\in\mathfrak{a}_{\theta}^{*}$. The $\phi$-critical exponent of $\Gamma$, denoted $\delta^{\phi}(\Gamma)\in[0,+\infty]$, is the critical exponent of the Poincaré series
$Q^{\phi}_{\Gamma}(s):=\sum_{\gamma\in\Gamma}e^{-s\phi(\kappa_{\theta}(\gamma))}=\sum_{\gamma\in\Gamma}e^{-s\phi(\kappa(\gamma))}.$
If $\delta^{\phi}(\Gamma)<+\infty$, a $\phi$-Patterson–Sullivan measure for $\Gamma$ (of dimension $\delta^{\phi}(\Gamma)$) is a probability measure on $\Lambda_{\theta}(\Gamma)$ such that for all $\gamma\in\Gamma$, $\gamma_{*}\mu$ and $\mu$ are mutually absolutely continuous and
$\frac{d\gamma_{*}\mu}{d\mu}(x)=e^{-\delta^{\phi}(\Gamma)\phi(B_{\theta}(\gamma^{-1},x))}$
for $\mu$-almost every $x\in\Lambda_{\theta}(\Gamma)$. By [12, Prop. 3.2], such measures exist.
In the case when $\Gamma$ is $\mathsf{P}_{\theta}$-transverse, notice that for any non-empty subset $\theta^{\prime}\subset\theta$, the forgetful projection $\operatorname{\mathcal{F}}_{\theta}\to\operatorname{\mathcal{F}}_{\theta^{\prime}}$ restricts to a homeomorphism between $\Lambda_{\theta}(\Gamma)$ and $\Lambda_{\theta^{\prime}}(\Gamma)$. Thus, for any $\phi\in\mathfrak{a}_{\theta}^{*}$ with $\delta^{\phi}(\Gamma)<+\infty$ and any $\phi$-Patterson–Sullivan measure $\mu$ for $\Gamma$ on $\Lambda_{\theta}(\Gamma)$, we may pushforward $\mu$ by this projection to obtain a measure on $\Lambda_{\theta^{\prime}}(\Gamma)$. We refer to this pushforward measure also as a $\phi$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{\theta^{\prime}}(\Gamma)$. Notice here that $\theta^{\prime}$ need not be symmetric, and $\phi$ need not lie in $\mathfrak{a}_{\theta^{\prime}}^{*}$.
Set
$\theta_{k}:=\{k-1,k,k+1,d-k-1,d-k,d-k+1\}\cap\{1,\dots,d-1\}.$
for every $k\in\{2,\dots,d-2\}$. In the remainder of the paper, when discussing Patterson–Sullivan measures, we will largely specialize to the case when $\Gamma$ is $\mathsf{P}_{\theta_{k}}$-Anosov and $\mu$ is a $\alpha_{k}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{k}(\Gamma)$. (Notice that $\alpha_{1}=2\omega_{1}-\omega_{2}$, $\alpha_{d-1}=2\omega_{d-1}-\omega_{d-2}$ and $\alpha_{k}=2\omega_{k}-\omega_{k-1}-\omega_{k+1}$ for every $k\in\{2,\dots,d-2\}$, so we may view $\alpha_{k}$ as an element in $\mathfrak{a}_{\theta_{k}}^{*}$ via the usual identification of $\mathfrak{a}_{\theta_{k}}^{*}$ as a subspace of $\mathfrak{a}^{*}$.) Since $\Gamma$ is $\mathsf{P}_{\theta_{k}}$-Anosov and $\alpha_{k}$ is positive on $\mathfrak{a}_{\theta_{k}}^{+}$, it follows from Sambarino [32, Cor. 5.4.3] that $\delta^{\alpha_{k}}(\Gamma)<+\infty$, and [32, Thm. A] that $\mu$ is unique.
2.5. Hyperconvex groups
With $\theta_{k}\subset\Delta$ as in the previous subsection, a subgroup $\Gamma\subset\mathsf{PSL}(d,\mathbb{K})$ is (dk)(d-k)-hyperconvex if it is $\mathsf{P}_{\theta_{k}}$-Anosov, and for all mutually distinct triples $x,y,z\in\Lambda_{\theta}(\Gamma)$ we have
$\Big((x^{k}\cap z^{d-k+1})+z^{d-k-1}\Big)\cap\Big((y^{k}\cap z^{d-k+1})+z^{d-k-1}\Big)=z^{d-k-1}.$
Notice that $\Gamma$ is (d1)(d-1)-hyperconvex if and only if for all mutually distinct triples $x,y,z\in\Lambda_{\theta_{1}}(\Gamma)$, we have
$x^{1}+y^{1}+z^{d-2}=\mathbb{K}^{d}.$
This condition was called $(1,1,2)$-hyperconvexity by Pozzetti–Sambarino–Wienhard [30]. They showed that in this setting the $\alpha_{1}$-Patterson–Sullivan measure behaves much like a conformal measure in the rank one setting. They also used this observation to bound the $\alpha_{1}$-critical exponent.
Theorem 2.3 ([30, Cor. 6.9 and 7.3]).
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (d1)(d-1)-hyperconvex, then the Hausdorff dimension of $\Lambda_{1}(\Gamma)$ is equal to $\delta^{\alpha_{1}}(\Gamma)$. Moreover, $\delta^{\alpha_{1}}(\Gamma)\leq 2$.
Farre–Pozzetti–Viaggi [19] showed that hyperconvex groups are always virtually isomorphic to convex cocompact Kleinian groups. We recall that two groups $G$ and $H$ are virtually isomorphic if there exists finite index subgroups $G^{\prime}\subset G$ and $H^{\prime}\subset H$ and finite normal subgroups $M\subset G^{\prime}$ and $N\subset H^{\prime}$ so that $G^{\prime}/M$ is isomorphic to $H^{\prime}/N$.
Theorem 2.4 ([19, Thm. A]).
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is $k$-hyperconvex for some $k\in\{1,\dots,d-1\}$, then $\Gamma$ is virtually isomorphic to a convex cocompact subgroup of $\mathsf{SL}(2,\mathbb{C})$.
They further adapt Ahlfors’ proof [2] that a convex cocompact Kleinian group is not cocompact, then its limit set has measure zero, to control the 2-dimensional Hausdorff measure of limit sets of hyperconvex groups.
Theorem 2.5 ([19, Thm. B]).
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex and $\partial\Gamma$ is not a two-sphere, then the 2-dimensional Hausdorff measure of $\Lambda_{k}(\Gamma)$ is zero.
We will often be able to reduce facts about (dk)(d-k)-hyperconvex groups to facts about (d1)(d-1)-hyperconvex groups using the exterior power representation. We explain this standard techniques more carefully in Appendix A, but we state the conclusion of this analysis here.
Proposition 2.6.
If $k\in\{2,\dots,d-1\}$, then there exist $d_{k}$, a representation $\eta_{k}:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\to\mathsf{PSL}(d_{k},\operatorname{\mathbb{K}})$ and a smooth embedding $E_{k}:\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\to\mathbb{\operatorname{\mathbb{P}}}(\operatorname{\mathbb{K}}^{d_{k}})$, which is holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$, such that if $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is (dk)(d-k)-hyperconvex subgroup then the following hold.
  1. (1)
    The subgroup $\eta_{k}(\Gamma)\subset\mathsf{PSL}(d_{k},\operatorname{\mathbb{K}})$ is $(d_{k}-1)$-hyperconvex, and
    $\Lambda_{1}(\eta_{k}(\Gamma))=E_{k}(\Lambda_{k}(\Gamma)).$
  2. (2)
    $\alpha_{1}(\kappa(\eta_{k}(g)))=\alpha_{k}(\kappa(g))$ for all $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ and hence $\delta^{\alpha_{1}}(\eta_{k}(\Gamma))=\delta^{\alpha_{k}}(\Gamma)=:\delta$.
  3. (3)
    If $\mu$ is the $\alpha_{k}$-Patterson-Sullivan measure on $\Lambda_{k}(\Gamma)$ for $\Gamma$ of dimension $\delta$, then $(E_{k})_{*}(\mu)$ is the $\alpha_{1}$-Patterson-Sullivan measure on $\Lambda_{1}(\eta(\Gamma))$ for $\eta_{k}(\Gamma)$ of dimension $\delta$.
(In the above, $\alpha_{1}$ is a root of $\mathsf{PSL}(d_{k},\operatorname{\mathbb{K}})$ while $\alpha_{k}$ is a root of $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$.)
As an example of how this is applied, notice that Theorem 2.3 and Proposition 2.6 have the following immediate corollary.
Corollary 2.7.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, then the Hausdorff dimension of $\Lambda_{k}(\Gamma)$ is equal to $\delta^{\alpha_{k}}(\Gamma)$. Moreover, $\delta^{\alpha_{k}}(\Gamma)\leq 2$.
2.6. Kleinian groups
Recall that $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ can be identified with $\mathsf{Isom}_{0}(\operatorname{\mathbb{H}}^{3})$, the connected component of the identity in the isometry group of real hyperbolic 3-space, in such a way that the geodesic boundary $\partial\operatorname{\mathbb{H}}^{3}$ identifies with $\mathbb{CP}^{1}$. Under this identification,
  1. (1)
    A subgroup $\Gamma\subset\mathsf{PSL}(2,\operatorname{\mathbb{C}})=\mathsf{Isom}_{0}(\operatorname{\mathbb{H}}^{3})$ is $\operatorname{\mathsf{P}}_{1}$-transverse if and only if it is discrete.
  2. (2)
    A subgroup $\Gamma\subset\mathsf{PSL}(2,\operatorname{\mathbb{C}})=\mathsf{Isom}_{0}(\operatorname{\mathbb{H}}^{3})$ is $\operatorname{\mathsf{P}}_{1}$-Anosov if and only if it is convex cocompact as a subgroup of $\mathsf{Isom}_{0}(\operatorname{\mathbb{H}}^{3})$.
We further note that the unique $\mathsf{PSU}(d)$-invariant probability measure on $\mathbb{CP}^{1}$ is a $\alpha_{1}$-Patterson–Sullivan measure of dimension two for any uniform lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
Remark 2.8.
There exist finitely generated Kleinian groups with critical exponent 2 whose limit set in $\mathbb{CP}^{1}$ has Hausdorff dimension two, but has Hausdorff measure zero (see [8]). Recall that all Kleinian groups are $\operatorname{\mathsf{P}}_{1}$-transverse.
3. Ahlfors regularity
Given a metric space $X$, a Borel measure $\mu$ on $X$ and $\delta>0$, we say that $\mu$ is $\delta$-Ahlfors regular if there are constants $D\geq 1$ and $t_{0}>0$ such that for every $x\in X$ and every $0<t<t_{0}$, we have
$\frac{1}{D}t^{\delta}\leq\mu(B(x,t))\leq Dt^{\delta},$
where $B(x,t)$ is the ball centered at $x$ of radius $t$. In this section we prove that if $\Gamma$ is (dk)(d-k)-hyperconvex, then its $\alpha_{k}$-Patterson–Sulivan measure is Ahlfors regular on $\Lambda_{k}(\Gamma)$.
Theorem 3.1.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, then the $\alpha_{k}$-Patterson-Sullivan measure $\mu$ for $\Gamma$ on $\Lambda_{k}(\Gamma)$ is $\delta^{\alpha_{k}}(\Gamma)$-Ahlfors regular. In particular, $\Lambda_{k}(\Gamma)$ has finite, non-zero $\delta^{\alpha_{k}}(\Gamma)$-dimensional Hausdorff measure.
Theorem 1.4 follows immediately. Also, as a consequence of Theorem 3.1 and results from Section 2.5, we may establish the first claim in Theorem 1.2.
Corollary 3.2.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex, then $\delta^{\alpha_{k}}(\Gamma)=2$ if and only if $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$.
Proof.
Suppose that $\delta^{\alpha_{k}}(\Gamma)=2$. By Theorem 3.1, the $2$-dimensional Hausdorff measure of $\Lambda_{k}(\Gamma)$ is positive, so Theorem 2.5 implies that $\partial\Gamma$ is a $2$-sphere. Then by Theorem 2.4, $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
Conversely, if $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$, then $\partial\Gamma$ is a two-sphere. Therefore, $\Lambda_{k}(\Gamma)$ has Hausdorff dimension at least two, so, by Theorem 2.3, $\delta^{\alpha_{1}}(\Gamma)\geq 2$. However, Theorem 2.3 also implies that $\delta^{\alpha_{1}}(\Gamma)\leq 2$, so $\delta^{\alpha_{1}}(\Gamma)=2$. ∎
3.1. Transverse representations and shadows
Let $\Omega\subset\mathbb{P}(\operatorname{\mathbb{R}}^{d_{0}})$ be a properly convex domain, i.e. is convex and bounded in some affine chart of $\mathbb{P}(\operatorname{\mathbb{R}}^{d_{0}})$. For any $a\in\partial\Omega$, a projective hyperplane $H\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{0}})$ is a supporting hyperplane to $\Omega$ through $a$ if $a\in H$ and $H\cap\Omega=\emptyset$. We say that $a\in\partial\Omega$ is a ${\mathcal{C}}^{1}$ point if there is a unique supporting hyperplane to $\Omega$ through $a$.
The Hilbert metric $\mathrm{d}_{\Omega}$ on $\Omega$ is defined as follows: Given $p,q\in\Omega$, let $a,b\in\partial\Omega$ be distinct points such that $a,p,q,b$ lie in this order along the projective line segment in $\Omega$ with $a$ and $b$ as endpoints. Then
$\mathrm{d}_{\Omega}(p,q):=\log\frac{\left|q-a\right|\left|p-b\right|}{\left|p-a\right|\left|q-b\right|},$
where $\left|\cdot\right|$ denotes some (any) norm on some (any) affine chart of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{0}})$ that contains $a$, $p$ ,$q$ and $b$. It is well-known that ${\rm d}_{\Omega}$ is an $\operatorname{\mathsf{Aut}}(\Omega)$-invariant, Finsler metric on $\Omega$, where $\operatorname{\mathsf{Aut}}(\Omega)$ denotes the subgroup of $\mathsf{PGL}({d_{0}},\operatorname{\mathbb{R}})$ that leaves $\Omega$ invariant.
Let $\Gamma_{0}\subset\operatorname{\mathsf{Aut}}(\Omega)$ be a discrete subgroup. The full orbital limit set of $\Gamma_{0}$ is
$\Lambda_{\Omega}(\Gamma_{0}):=\left\{x\in\partial\Omega\ |\ x=\lim\gamma_{n}(p)\ \text{ for some }p\in\Omega\text{ and some }\{\gamma_{n}\}\subset\Gamma_{0}\right\}.$
We may specify a collection of open subsets of $\Lambda_{\Omega}(\Gamma_{0})$ called shadows in the following way. For any $p,q\in\Omega$ and $r>0$, let $\mathcal{O}_{r}(p,q)$ denote the set of points $a\in\partial\Omega$ for which the projective line segment $[p,a)$ in $\Omega$ between $p$ and $a$ intersects the open ball $B_{\Omega}(q,r)$ of radius $r$ (in the Hilbert metric on $\Omega$) centered at $q$. We refer to
$\widehat{\mathcal{O}}_{r}(p,q):=\mathcal{O}_{r}(p,q)\cap\Lambda_{\Omega}(\Gamma_{0})$
as the $r$-shadow in $\Lambda_{\Omega}(\Gamma_{0})$ from $p$ centered at $q$.
Given a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{0}})$, we say that a discrete subgroup $\Gamma_{0}\subset\operatorname{\mathsf{Aut}}(\Omega)$ is projectively visible if every point in $\Lambda_{\Omega}(\Gamma_{0})$ is a ${\mathcal{C}}^{1}$ point, and any two distinct points in $\Lambda_{\Omega}(\Gamma_{0})$ are joined by a projective line segment in $\Omega$. For any such $\Gamma_{0}$, a representation $\rho:\Gamma_{0}\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is called $\mathsf{P}_{\theta}$-transverse if there exists a continuous embedding
$\xi:\Lambda_{\Omega}(\Gamma_{0})\rightarrow\mathcal{F}_{\theta}(\operatorname{\mathbb{K}}^{d})$
with the following properties:
  1. (1)
    $\xi$ is $\rho$-equivariant, i.e. $\rho(\gamma)\circ\xi=\xi\circ\gamma$ for all $\gamma\in\Gamma$,
  2. (2)
    $\xi$ is transverse, i.e. $\xi(\Lambda_{\Omega}(\Gamma))$ is a transverse subset of $\mathcal{F}_{\theta}(\operatorname{\mathbb{K}}^{d})$,
  3. (3)
    $\xi$ is strongly dynamics preserving, i.e. if $(\gamma_{i})_{i=1}^{\infty}$ is a sequence in $\Gamma$ so that $\gamma_{i}(p)\to a\in\Lambda_{\Omega}(\Gamma)$ and $\gamma_{i}^{-1}(p)\to b\in\Lambda_{\Omega}(\Gamma)$ for some (any) $p\in\Omega$, then $\rho(\gamma_{i})(x)\to\xi(a)$ if $x\in\mathcal{F}_{\theta}(\operatorname{\mathbb{K}}^{d})$ is transverse to $\xi(b)$.
We refer to $\xi$ as the limit map of $\rho$ (note that this is unique to $\rho$).
It follows from Proposition 2.1 that if $\rho:\Gamma_{0}\to\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{\theta}$-transverse representation with limit map $\xi$, then $\rho(\Gamma_{0})$ is a $\mathsf{P}_{\theta}$-transverse group with $\xi(\Lambda_{\Omega}(\Gamma_{0}))$ as its $\theta$-limit set. The following theorem states that the converse is also true when $\theta=\{k,d-k\}$.
Theorem 3.3 ([12, Thm. 6.2],[38, Cor. 1.30]).
If $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{k,d-k}$-transverse subgroup for some $k\leq d-k$, then there exist some $d_{0}\in\operatorname{\mathbb{N}}$, a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{0}})$, a projectively visible subgroup $\Gamma_{0}\subset\operatorname{\mathsf{Aut}}(\Omega)$, and a faithful $\mathsf{P}_{k,d-k}$-transverse representation $\rho:\Gamma_{0}\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ with limit map $\xi:\Lambda_{\Omega}(\Gamma_{0})\to\mathcal{F}_{k,d-k}$ so that
  1. (1)
    $\rho(\Gamma_{0})=\Gamma$,
  2. (2)
    $\xi(\Lambda_{\Omega}(\Gamma_{0}))=\Lambda_{k,d-k}(\Gamma)$,
  3. (3)
    $\alpha_{1}(\kappa(\gamma))=\alpha_{k}(\kappa(\rho(\gamma)))$ for all $\gamma\in\Gamma_{0}$.
Furthermore, if $\Gamma$ is $\operatorname{\mathsf{P}}_{k,d-k}$-Anosov, then $\Gamma_{0}$ acts cocompactly on the convex hull in $\Omega$ of $\Lambda_{\Omega}(\Gamma_{0})$.
If $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{1,d-1}$-transverse subgroup, then by making a choice of $d_{0}$, $\Omega$, $\Gamma_{0}$, and $\rho$ as given by Theorem 3.3 (with $k=1$), we may define shadows in $\Lambda_{1}(\Gamma)$ as the images of the shadows in $\Lambda_{\Omega}(\Gamma_{0})$ under $\xi^{1}$, where $\xi=(\xi^{1},\xi^{d-1})$ is the limit map of $\rho$. In our previous work [12, Prop. 7.1], we proved a shadow lemma for these shadows, i.e. we estimated the $\phi$-Patterson–Sullivan measures of the shadows in $\Lambda_{1}(\Gamma)$. Also, assuming an analog of the (d1)(d-1)-hyperconvexity assumption for transverse representations, we estimated the in-radii and out-radii of these shadows [11, Thm. 5.1 and 7.1]. We will now state a specialization of both of these results to the setting where $\Gamma$ is (d1)(d-1)-hyperconvex and $\phi=\alpha_{1}$.
Theorem 3.4 ([12, Prop. 7.1], [11, Thm. 5.1 and 7.1]).
Let $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ be a (d1)(d-1)-hyperconvex subgroup and let $\mu$ be the $\alpha_{1}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{1}(\Gamma)$. Let $d_{0}\in\operatorname{\mathbb{N}}$, $\Omega\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{0}})$ be a properly convex domain, $\Gamma_{0}\subset\operatorname{\mathsf{Aut}}(\Omega)$ be a projectively visible group, and $\rho:\Gamma_{0}\to\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ be a faithful, $\mathsf{P}_{1,d-1}$-transverse representation with limit map $\xi:\Lambda_{\Omega}(\Gamma_{0})\to\operatorname{\mathcal{F}}_{1,d-1}$ satisfying properties (1)–(3) of Theorem 3.3 (with $k=1$). Fix a base point $o\in\Omega$. There exists $r_{0}>0$ such that for all $r>r_{0}$, there exists $C(r)>1$ for which the following hold:
  1. (i)
    For all $\gamma\in\Gamma_{0}$,
    $\frac{1}{C(r)}e^{-\delta\alpha_{1}(\kappa(\rho(\gamma)))}\leq\mu\Big(\xi^{1}\big(\widehat{\mathcal{O}}_{r}(o,\gamma(o))\big)\Big)\leq C(r)e^{-\delta\alpha_{1}(\kappa(\rho(\gamma)))}.$
  2. (ii)
    If $a\in\Lambda_{\Omega}(\Gamma_{0})$, $p$ lies in the projective line segment $[o,a)$ in $\Omega$ with endpoints $o$ and $a$, and $\gamma\in\Gamma_{0}$ satisfies ${\rm d}_{\Omega}(p,\gamma(o))<r$, then
    $B\left(\xi^{1}(a),\frac{1}{C(r)}e^{-\alpha_{1}(\kappa(\rho(\gamma)))}\right)\cap\Lambda_{1}(\Gamma)\subset\xi^{1}\left(\widehat{\mathcal{O}}_{r}(o,p)\right)\subset B\left(\xi^{1}(a),C(r)e^{-\alpha_{1}(\kappa(\rho(\gamma)))}\right)\cap\Lambda_{1}(\Gamma),$
    where $B(x,t)$ denotes the ball centered at $x\in\operatorname{\mathcal{F}}_{1}(\operatorname{\mathbb{K}}^{d})$ of radius $t$ in the Riemannian metric on $\operatorname{\mathcal{F}}_{1}(\operatorname{\mathbb{K}}^{d})$.
The results in [11] used closed shadows, but due to the nesting properties of these sets, the difference is immaterial.
3.2. Proof of Theorem 3.1
Using Proposition 2.6 it suffices to consider the case when $k=1$.
Suppose that $\Gamma\subset\mathsf{PSL}(d,\mathbb{K})$ is (d1)(d-1)-hyperconvex and $\mu$ is the $\alpha_{1}$-Patterson–Sullivan measure for $\Gamma$. By Theorem 3.3, there exist $d_{0}\in\operatorname{\mathbb{N}}$, a properly convex domain $\Omega\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{R}}^{d_{0}})$, a projectively visible subgroup $\Gamma_{0}\subset\operatorname{\mathsf{Aut}}(\Omega)$, and a faithful, $\mathsf{P}_{1,d-1}$-transverse representation $\rho:\Gamma_{0}\to\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ with image $\Gamma$ and limit map $\xi:\Lambda_{\Omega}(\Gamma_{0})\to\Lambda_{1,d-1}(\Gamma)$ that satisfy properties (1)–(3) of Theorem 3.3 (with $k=1$). Since $\Gamma$ is $\mathsf{P}_{1,d-1}$-Anosov, $\Gamma_{0}$ acts cocompactly on the convex hull $\mathcal{C}$ in $\Omega$ of $\Lambda_{\Omega}(\Gamma_{0})$. So, if we choose $o\in\mathcal{C}$, then there exists $R>0$ so that if $p\in\mathcal{C}$ then there exists $\gamma\in\Gamma_{0}$ so that $d_{\Omega}(p,\gamma(o))<R$ (in the Hilbert metric on $\Omega$). We may also assume that $R>r_{0}$, where $r_{0}>0$ is the constant in Theorem 3.4.
Fix $x\in\Lambda_{1}(\Gamma)$. Then there exists $a\in\Lambda_{\Omega}(\Gamma_{0})$ such that $\xi^{1}(a)=x$. Fix a base point $o\in\Omega$, and let $[o,a)$ denote the projective line segment in $\Omega$ with endpoints $o$ and $a$. Recall that ${\rm d}_{\Omega}$ denotes the Hilbert metric on $\Omega$ and for any $p\in\Omega$ and $r>0$, $B_{\Omega}(p,r)$ denotes the open Hilbert ball of radius $r$ centered at $p$.
Iteratively construct a sequence $\{\gamma_{n}\}$ in $\Gamma_{0}$ as follows. Set $\gamma_{1}:=\operatorname{\mathrm{id}}$. Then given $\gamma_{n}$, choose $\gamma_{n+1}\in\{\gamma\in\Gamma_{0}:B_{\Omega}(\gamma(o),R)\cap[o,a)\neq\emptyset\}$ such that the endpoint of the interval $B_{\Omega}(\gamma_{n}(o),R)\cap[o,a)$ that is furthest from $o$ lies in $B_{\Omega}(\gamma_{n+1}(o),R)$. Observe that ${\rm d}_{\Omega}(\gamma_{n}(o),\gamma_{n+1}(o))<2R$ for all integers $n\geq 1$, and $o,a)\subset\bigcup_{n=1}^{\infty}B_{\Omega}(\gamma_{n}(o),R)$.
For each positive integer $n$, set $\beta_{n}:=\gamma_{n}^{-1}\gamma_{n+1}$. It is a standard linear algebra fact that for all $g_{1},g_{2}\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ and all $i\in\{1,\dots,d\}$, we have
$\sigma_{i}(g_{1})\sigma_{d}(g_{2})\leq\sigma_{i}(g_{1}g_{2})\leq\sigma_{i}(g_{1})\sigma_{1}(g_{2}).$
Thus, for all positive integers $n$, we have
$\alpha_{1}(\kappa(\rho(\gamma_{n})))-\log\frac{\sigma_{1}(\rho(\beta_{n}))}{\sigma_{d}(\rho(\beta_{n}))}\leq\alpha_{1}(\kappa(\rho(\gamma_{n+1})))\leq\alpha_{1}(\kappa(\rho(\gamma_{n})))+\log\frac{\sigma_{1}(\rho(\beta_{n}))}{\sigma_{d}(\rho(\beta_{n}))}.$
Since ${\rm d}_{\Omega}(b,\beta_{n}(b))\leq 2R$ for all $n$, if we set
$S:=\max\left\{\frac{\sigma_{1}(\rho(\beta))}{\sigma_{d}(\rho(\beta))}:\beta\in\Gamma_{0}\text{ and }{\rm d}_{\Omega}(b,\beta(b))\leq 2R\right\},$
then for all $n$, we have
$\left|\alpha_{1}(\kappa(\rho(\gamma_{n})))-\alpha_{1}(\kappa(\rho(\gamma_{n+1})))\right|\leq\log S.$ (1)
For each $r>r_{0}$, let $C(r)\geq 1$ be the constant given by Theorem 3.4. Set $\delta:=\delta^{\alpha_{1}}(\Gamma)$, $t_{0}:=\frac{1}{SC(R)}$, and $D:=\max\{C(2R)C(R)^{\delta}S^{\delta},C(R)C(2R)^{\delta}S^{\delta}\}$. To prove Ahlfors regularity, it suffices to show that for all $0<t<t_{0}$, we have
$\frac{1}{D}t^{\delta}\leq\mu(B(x,t))\leq Dt^{\delta}.$ (2)
First, we prove the required upper bound for $\mu(B(x,t))$. Since $\rho$ is $\mathsf{P}_{1,d-1}$-divergent and $\{\gamma_{n}\}$ is a diverging sequence in $\Gamma_{0}$, we have $\alpha_{1}(\kappa(\rho(\gamma_{n})))\to\infty$, so we may let $\ell$ be the smallest positive integer such that $\alpha_{1}(\kappa(\rho(\gamma_{\ell})))\geq-\log(SC(R)t)$. Since $\gamma_{1}=\operatorname{\mathrm{id}}$ and $0<-\log(SC(R)t)$, observe that $\ell\geq 2$ and $\alpha_{1}(\kappa(\rho(\gamma_{\ell-1})))<-\log(SC(R)t)$. It now follows from (1) that
$\log(SC(R)t)\leq\alpha_{1}(\kappa(\rho(\gamma_{\ell})))\leq-\log(SC(R)t)+\log S=-\log(C(R)t),$
or equivalently, that
$\frac{1}{SC(R)}e^{-\alpha_{1}(\kappa(\rho(\gamma_{\ell})))}\leq t\leq\frac{1}{C(R)}e^{-\alpha_{1}(\kappa(\rho(\gamma_{\ell})))}.$ (3)
Choose $p\in[o,a)$ so that ${\rm d}_{\Omega}(p,\gamma_{\ell}(o))<R$. By the first inclusion of Theorem 3.4 part (ii) and the second inequality in (3), we have that
$B(x,t)\cap\Lambda_{1}(\Gamma)\subset\xi^{1}(\widehat{\mathcal{O}}_{R}(o,a))\subset\xi^{1}(\widehat{\mathcal{O}}_{2R}(o,\gamma_{\ell}(o))).$
Then the second inequality of Theorem 3.4 part (i) and the first inequality in (3) imply that
$\mu(B(x,t))\leq\mu\left(\xi^{1}(\widehat{\mathcal{O}}_{2R}(o,\gamma_{\ell}(o)))\right)\leq C(2R)e^{-\delta\alpha_{1}(\kappa(\rho(\gamma_{\ell})))}\leq C(2R)(SC(R))^{\delta}t^{\delta}\leq Dt^{\delta}.$
The proof of the required lower bound for $\mu(B(x,t))$ is similar. Observe that since $0<t<SC(2R)$ (because $t_{0}<1\leq SC(2R)$) and $\gamma_{1}=\operatorname{\mathrm{id}}$, we may let $m$ be the largest positive integer such that $\alpha_{1}(\kappa(\rho(\gamma_{m})))\leq-\log\frac{t}{SC(2R)}$. Then $\alpha_{1}(\kappa(\rho(\gamma_{m+1})))\geq-\log\frac{t}{SC(2R)}$, so (1) implies that
$\log\frac{t}{C(2R)}=-\log\frac{t}{SC(2R)}-\log S\leq\alpha_{1}(\kappa(\rho(\gamma_{m})))\leq-\log\frac{t}{SC(2R)},$
or equivalently that
$C(2R)e^{-\alpha_{1}(\kappa(\rho(\gamma_{m})))}\leq t\leq SC(2R)e^{-\alpha_{1}(\kappa(\rho(\gamma_{m})))}.$ (4)
Choose $p\in[o,a)$ so that ${\rm d}_{\Omega}(p,\gamma_{m}(o))<R$. By the second inclusion of Theorem 3.4 part (ii) and the first inequality of (4),
$\xi^{1}\Big(\widehat{\mathcal{O}}_{R}(o,\gamma_{m}(o))\Big)\subset\xi^{1}\Big(\widehat{\mathcal{O}}_{2R}(o,p)\Big)\subset B(x,t)\cap\Lambda_{1}(\Gamma).$
Then the first inequality of Theorem 3.4 part (i) and the second inequality in (4) imply that
$\mu\Big(B(x,t)\Big)\geq\mu\left(\xi^{1}(\widehat{\mathcal{O}}_{R}(o,\gamma_{m}(o)))\right)\geq\frac{1}{C(R)}e^{-\delta\alpha_{1}(\kappa(\rho(\gamma_{m})))}\geq\frac{t^{\delta}}{C(R)(C(2R)S)^{\delta}}\geq\frac{1}{D}t^{\delta}.$
This completes the proof of Equation (2) and hence the proof that $\mu$ is $\delta$-Ahlfors regular.
We now recall the standard argument that the support of a $\delta$-Ahlfors regular measure has finite non-zero $\delta$-dimensional Hausdorff measure. First, suppose that $\{B(x_{i},t_{i})\}$ is a countable cover of $\Lambda_{1}(\Gamma)$ by balls of radius $t_{i}<t_{0}$ based at points $x_{i}\in\Lambda_{1}(\Gamma).$ Then, by Equation (2) and the fact that $\mu$ is a probability measure, we have
$\sum_{i=1}^{\infty}t_{i}^{\delta}\geq\frac{1}{D}\sum_{i=1}^{\infty}\mu(B(x_{i},t_{i}))\geq\frac{1}{D}.$
It follows that the $\delta$-dimensional Hausdorff measure of $\Lambda_{1}(\Gamma)$ is at least $\frac{1}{D}$, hence non-zero.
On the other hand, if $0<t<t_{0}$, the Vitali covering lemma implies that we may find a finite cover $\{B(x_{i},t)\}_{i=1}^{k}$ of $\Lambda_{1}(\Gamma)$ by balls of radius $t$ so that the collection $\{B(x_{i},\frac{t}{3})\}_{i=1}^{k}$ is a disjoint collection of $\frac{t}{3}$-balls. So
$1\geq\sum_{i=1}^{k}\mu\Big(B(x_{i},\frac{t}{3})\Big)\geq\frac{kt^{\delta}}{D3^{\delta}}$
which implies that $\delta$-dimensional Hausdorff measure of $\Lambda_{1}(\Gamma)$ is at most $D3^{\delta}$, hence finite. ∎
4. Simplicity of the Zariski closure
Notice that $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is the identity component of $\mathsf{PGL}(d,\operatorname{\mathbb{K}})$, which is a real algebraic group. Hence, given a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, its real Zariski closure in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is the intersection of $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ with the real Zariski closure of $\Gamma$ in $\mathsf{PGL}(d,\operatorname{\mathbb{K}})$. We also say that a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is strongly irreducible if every finite index subgroup of $\Gamma$ is irreducible. Notice that if $\operatorname{\mathsf{G}}^{0}$ is the identity component of the real Zariski closure of a discrete subgroup $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, then $\Gamma^{0}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$ is a finite index subgroup of $\Gamma$. Thus, if $\Gamma$ is strongly irreducible, then $\Gamma^{0}$, and hence $\operatorname{\mathsf{G}}^{0}$, is irreducible.
In this section we establish a sufficient condition for the real Zariski closure of a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup to be simple.
Theorem 4.1.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and when $k>1$ assume that $\Gamma$ is also $\operatorname{\mathsf{P}}_{k-1}$-Anosov. If $\Lambda_{k}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$, then the identity component of the real Zariski closure of $\Gamma$ is a simple Lie group with trivial center.
Example 4.2.
When $k>1$, there exist strongly irreducible $\mathsf{P}_{k}$-Anosov subgroups where the limit set is a ${\mathcal{C}}^{1}$-submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$, but the real Zariski closure of $\Gamma$ is not simple. For instance, suppose $\Gamma_{0}\subset\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is a uniform lattice, and let $\Gamma_{0}^{\prime}\subset\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{C}})$ denotes its lift to $\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{C}})$. By replacing $\Gamma_{0}^{\prime}$ by a finite index subgroup we can assume that $\Gamma_{0}^{\prime}$ admits a surjective homomorphism to a non-abelian free group, see Agol–Groves–Manning [1]. Then, since non-abelian free groups admit Zariski dense representations into $\mathsf{SU}(k)$, we can construct a Zariski dense representation $\tau_{0}^{\prime}:\Gamma_{0}\rightarrow\mathsf{SU}(k)$.
Next let $\tau^{\prime}:\Gamma_{0}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(\operatorname{\mathbb{C}}^{2}\otimes\operatorname{\mathbb{C}}^{k})$ be the representation defined by
$\tau(\gamma)(v\otimes w)=(\gamma(v))\otimes(\tau_{0}(\gamma)(w)),$
and let $\tau:=\pi\circ\tau^{\prime}$, where $\pi:\operatorname{\mathsf{SL}}(\operatorname{\mathbb{C}}^{2}\otimes\operatorname{\mathbb{C}}^{k})\to\mathsf{PSL}(\operatorname{\mathbb{C}}^{2}\otimes\operatorname{\mathbb{C}}^{k})$ is the quotient map. Using the fact that $\Gamma_{0}<\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ is $\operatorname{\mathsf{P}}_{1}$-Anosov and has limit set $\Lambda_{1}(\Gamma)=\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{2})$, one can check that $\Gamma:=\tau(\Gamma_{0}^{\prime})$ is $\mathsf{P}_{k}$-Anosov with limit set
$\Lambda_{k}(\Gamma)=\{L\otimes\operatorname{\mathbb{C}}^{k}:L\in\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{2})\}.$
Then $\Lambda_{k}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold and the real Zariski closure of $\Gamma$ is isomorphic to
$\mathsf{P}(\operatorname{\mathsf{SL}}(2,\operatorname{\mathbb{C}})\oplus\mathsf{SU}(k)).$
4.1. Anosov groups in general semisimple Lie groups
A Lie group is reductive if its Lie algebra decomposes as a direct sum of simple Lie algebras and an Abelian Lie algebra (which is necessarily the center of its Lie algebra). Equivalently, it admits a representation that is completely reducible, i.e. decomposes into a direct sum of irreducible representations. A reductive Lie group is semisimple if its Lie algebra has trivial center.
If a Lie group $\operatorname{\mathsf{G}}$ is semisimple, then its Lie algebra $\mathfrak{g}$ admits a Cartan decomposition, i.e. we may write
$\mathfrak{g}=\mathfrak{k}\oplus\mathfrak{p},$
where $\mathfrak{k}\subset\mathfrak{g}$ is a maximal compact Lie subalgebra and $\mathfrak{p}$ is the orthogonal complement of $\mathfrak{k}$ with respect to the Killing form. Let $\mathfrak{a}_{\operatorname{\mathsf{G}}}\subset\mathfrak{p}$ be a maximal Abelian subspace. For every linear map $\alpha\in\mathfrak{a}_{\operatorname{\mathsf{G}}}^{*}$, let
$\mathfrak{g}_{\alpha}:=\{X\in\mathfrak{a}_{\operatorname{\mathsf{G}}}:[X,Y]=\alpha(X)Y\text{ for all }Y\in\mathfrak{a}_{\operatorname{\mathsf{G}}}\}.$
We say that $\alpha\in\mathfrak{a}_{\operatorname{\mathsf{G}}}^{*}$ is a restricted root of $\operatorname{\mathsf{G}}$ if $\mathfrak{g}_{\alpha}$ is non-zero. Let $\Phi$ denote the set of restricted simple roots of $\operatorname{\mathsf{G}}$. Then we have the restricted root space decomposition
$\mathfrak{g}=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Phi}\mathfrak{g}_{\alpha}.$
Choose a vector $X_{0}\in\mathfrak{a}_{\operatorname{\mathsf{G}}}$ such that $\alpha(X_{0})\neq 0$ for all $\alpha\in\Phi$. Then we may decompose $\Phi=\Phi_{-}\cup\Phi_{+}$, where $\Phi_{\pm}:=\{\alpha\in\Phi:{\rm sgn}(\alpha(X_{0}))=\pm 1\}$. It is a fact that $\Phi_{-}=-\Phi_{+}$, and that there is a basis $\Delta_{\operatorname{\mathsf{G}}}\subset\Phi_{+}$ of $\mathfrak{a}_{\operatorname{\mathsf{G}}}^{*}$ such that every linear map in $\Phi^{+}$ can be written as a non-negative linear combination of the elements in $\Delta_{\operatorname{\mathsf{G}}}$. The linear maps in $\Delta_{\operatorname{\mathsf{G}}}$ are called the restricted simple roots of $\operatorname{\mathsf{G}}$.
Given a non-empty subset $\theta\subset\Delta_{\operatorname{\mathsf{G}}}$, let
$\mathfrak{u}_{\theta}:=\bigoplus_{\alpha\in\Phi\setminus\operatorname{Span}_{\operatorname{\mathbb{R}}}(\Delta_{\operatorname{\mathsf{G}}}\setminus\theta)}\mathfrak{g}_{\alpha}.$
The $\theta$-standard parabolic subgroup is the subgroup $\operatorname{\mathsf{P}}_{\theta}\subset\operatorname{\mathsf{G}}$ that normalizes $\mathfrak{u}_{\theta}$, and the $\theta$-flag manifold is the compact manifold $\operatorname{\mathcal{F}}_{\theta}:=\operatorname{\mathsf{G}}/\operatorname{\mathsf{P}}_{\theta}$. Note that the left action of $\operatorname{\mathsf{G}}$ on $\operatorname{\mathcal{F}}_{\theta}$ endows $\operatorname{\mathcal{F}}_{\theta}$ with the structure of a homogeneous $\operatorname{\mathsf{G}}$-space.
The (closed) positive Weyl chamber is
$\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}:=\{X\in\mathfrak{a}:\alpha(X)\geq 0\text{ for all }\alpha\in\Delta_{\operatorname{\mathsf{G}}}\}.$
The $\mathsf{KAK}$-decomposition theorem states that every $g\in\operatorname{\mathsf{G}}$ can be written as $g=k_{1}e^{\kappa(g)}k_{2}$ for unique $\kappa(g)\in\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}$ and some $k_{1}$ and $k_{2}$ in $\mathsf{K}$, which is the maximal compact subgroup of $\operatorname{\mathsf{G}}$ whose Lie algebra is $\mathfrak{k}$. This defines for us the Cartan projection
$\kappa:\operatorname{\mathsf{G}}\to\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}.$
In the case when $\operatorname{\mathsf{G}}=\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, we may choose $\mathfrak{k}$ to be $\mathfrak{su}(d,\operatorname{\mathbb{K}})$, which is the set of traceless, skew-Hermitian, $d\times d$, $\operatorname{\mathbb{K}}$-valued matrices, and choose $\mathfrak{a}_{\operatorname{\mathsf{G}}}$ and $\mathfrak{a}_{\operatorname{\mathsf{G}}}^{+}$ to be $\mathfrak{a}$ and $\mathfrak{a}^{+}$ as defined in Section 2.1. One can then verify that the restricted simple roots defined here agree with the ones defined in Section 2.1, and that $\Delta_{\operatorname{\mathsf{G}}}=\{\alpha_{k}:k\in\Delta\}$. Also, the Cartan projection defined here agrees with the one defined in Section 2.1.
Let $\theta\subset\Delta_{\operatorname{\mathsf{G}}}$ be a non-empty, symmetric subset. A discrete subgroup $\Gamma\subset\operatorname{\mathsf{G}}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov if it is finitely generated, and there exists $a,C>0$ such that
$\alpha_{k}(\kappa(\gamma))\geq a|\gamma|-C$
for all $\gamma\in\Gamma$, and $k\in\theta$, where $|\gamma|$ is the word length $\gamma$ with respect to some fixed finite generating set for $\Gamma$. By Theorem 2.2, in the case when $\operatorname{\mathsf{G}}=\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, this agrees with the notion of $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroups defined in Section 2.3. In general, every $\operatorname{\mathsf{P}}_{\theta}$-Anosov subgroup $\Gamma\subset\operatorname{\mathsf{G}}$ admits a unique non-empty, closed, $\Gamma$-invariant, subset $\Lambda_{\theta}(\Gamma)\subset\operatorname{\mathcal{F}}_{\theta}$ whose points are pairwise transverse, and on which the $\Gamma$-action is strongly dynamics preserving, i.e. if $\{\gamma_{n}\}$ is an infinite sequence in $\Gamma$, then up to taking a subsequence, there exist $a,b\in\Lambda_{\theta}(\Gamma)$ such that the sequence $\{\gamma_{n}\}$, when restricted to the set of flags in $\operatorname{\mathcal{F}}_{\theta}$ that are transverse to $b$, converges uniformly on compact sets to the constant map whose image is $a$. We refer to $\Lambda_{\theta}(\Gamma)$ as the $\theta$-limit set of $\Gamma$.
Results in [20] imply the following.
Lemma 4.3.
Suppose that $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a $\mathsf{P}_{k}$-Anosov subgroup that is strongly irreducible, let $\mathsf{G}$ be the real Zariski closure of $\Gamma$ in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, and let $\operatorname{\mathsf{G}}^{0}$ denote the identity component of $\operatorname{\mathsf{G}}$. Then:
  1. (1)
    $\operatorname{\mathsf{G}}^{0}$ has trivial center and is semisimple.
  2. (2)
    There exists a non-empty $\theta\subset\Delta_{\operatorname{\mathsf{G}}}$ such that $\Gamma$ is $\mathsf{P}_{\theta}$-Anosov, and for which there is a $\mathsf{G}$-equivariant smooth embedding $\iota:\mathsf{G}/\mathsf{P}_{\theta}\hookrightarrow\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ such that the $\theta$-limit set $\Lambda_{\theta}(\Gamma)\subset\mathsf{G}/\mathsf{P}_{\theta}$ of $\Gamma$ in satisfies
    $\iota(\Lambda_{\theta}(\Gamma))=\Lambda_{k}(\Gamma).$
Proof.
Let $\Gamma^{\prime}$ and $\operatorname{\mathsf{G}}^{\prime}$ respectively denote the preimages of $\Gamma$ and $\operatorname{\mathsf{G}}$ under the quotient map $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\to\mathsf{PSL}(d,\operatorname{\mathbb{K}})$. Observe that $\operatorname{\mathsf{G}}^{\prime}$ is the real Zariski closure of $\Gamma^{\prime}$.
First, we prove (1). Since $\Gamma$ is strongly irreducible, it follows that the identity component $(\operatorname{\mathsf{G}}^{\prime})^{0}$ acts irreducibly on $\operatorname{\mathbb{K}}^{d}$, and hence is reductive. We may then apply Schur’s lemma to deduce that the center of $(\operatorname{\mathsf{G}}^{\prime})^{0}$ acts on $\operatorname{\mathbb{K}}^{d}$ by scaling, and hence lies in the center of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$. In particular, $(\operatorname{\mathsf{G}}^{\prime})^{0}$ is semisimple. It now follows that $\operatorname{\mathsf{G}}^{0}=\pi((\operatorname{\mathsf{G}}^{\prime})^{0})$ is semisimple and has trivial center.
To prove (2), notice that it suffices to prove the statement for $\operatorname{\mathsf{G}}^{\prime}$ and $\Gamma^{\prime}$ in place of $\operatorname{\mathsf{G}}$ and $\Gamma$. Let $\eta=\eta_{k}:\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\to\operatorname{\mathsf{SL}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ be the representation induced by the natural linear action of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ on $\wedge^{k}\operatorname{\mathbb{K}}^{d}$, and let $\chi:\mathfrak{a}_{\operatorname{\mathsf{G}}^{\prime}}\to\operatorname{\mathbb{R}}$ be the highest weight of $\eta|_{\operatorname{\mathsf{G}}^{\prime}}$. We may write
$\chi=\sum_{\alpha\in\Delta_{\operatorname{\mathsf{G}}^{\prime}}}c_{\alpha}\omega_{\alpha},$
where $\omega_{\alpha}$ is the restricted simple weight corresponding to $\alpha$ and each $c_{\alpha}$ is a non-negative integer. Let $\theta:=\{\alpha\in\Delta:c_{\alpha}>0\}$. Since $\Gamma$ is $\operatorname{\mathsf{P}}_{k}$-Anosov, $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{1}$-Anosov (see Item (VII) in Appendix 2.6), and so contains a proximal element. In the language of Guichard–Gueritaud–Kassel–Wienhard (see [20, Lemma 3.2]), we have just observed that $\eta|_{\operatorname{\mathsf{G}}^{\prime}}$ is $\theta$-proximal, so we may apply [20, Propositions 3.3 and 3.5] to deduce (2). ∎
Given Lemma 4.3 part (1), to prove Theorem 4.1, it now suffices to show that the real Zariski closure of $\operatorname{\mathsf{G}}$ in $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is simple. The proof of this is divided into two steps which are written as Lemmas 4.5 and 4.9 below. The former shows that the real Zariski closure of $\Gamma$ has no compact factors, while the latter shows that it has at most one simple factor.
4.2. Ruling out compact simple factors
We prove some general results which show that if $\Gamma$ satisfies the hypothesis Theorem 4.1, then the connected component of the identity of the real Zariski closure of $\Gamma$ has no compact simple factors.
We start by recording a standard fact about irreducible representations of products of semisimple Lie groups.
Lemma 4.4.
Suppose $\mathsf{H}_{1},\mathsf{H}_{2}\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ are commuting connected reductive Lie groups and $\mathsf{H}:=\mathsf{H}_{1}\mathsf{H}_{2}$ acts irreducibly on $\operatorname{\mathbb{K}}^{d}$. Then there exist irreducible representations $\rho_{j}:\mathsf{H}_{j}\rightarrow\operatorname{\mathsf{SL}}(d_{j},\operatorname{\mathbb{K}})$ and a linear identification
$\operatorname{\mathbb{K}}^{d}=\operatorname{\mathbb{K}}^{d_{1}}\otimes\operatorname{\mathbb{K}}^{d_{2}}$
such that elements $h=h_{1}h_{2}\in\mathsf{H}$ act by
$h(v_{1}\otimes v_{2})=(\rho_{1}(h_{1})(v_{1}))\otimes(\rho_{2}(h_{2})(v_{2})).$
Proof.
Since $\mathsf{H}_{1}$ is reductive, we may decompose its action on $\operatorname{\mathbb{K}}^{d}$ into its irreducible factors, i.e. there exist $m_{1},\dots,m_{k},n_{1},\dots,n_{k}\in\operatorname{\mathbb{N}}$ and irreducible representations $\tau_{j}:\mathsf{H}_{1}\rightarrow\operatorname{\mathsf{SL}}(m_{j},\operatorname{\mathbb{K}})$ so that we can identify
$\operatorname{\mathbb{K}}^{d}=\oplus_{j=1}^{k}(\operatorname{\mathbb{K}}^{m_{j}})^{n_{j}}$
and $\mathsf{H}_{1}$ acts by $(\tau_{1}^{n_{1}},\dots,\tau_{k}^{n_{k}})$. Since $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ commute, $\mathsf{H}_{2}$ preserves each $(\operatorname{\mathbb{K}}^{m_{j}})^{n_{j}}\subset\operatorname{\mathbb{K}}^{d}$. Then, since $\mathsf{H}$ acts irreducibly, we must have $k=1$.
Relative to the decomposition $\operatorname{\mathbb{K}}^{d}=(\operatorname{\mathbb{K}}^{m_{1}})^{n_{1}}$, we can write each element $h\in\mathsf{H}_{2}$ as
$h=(T_{h,u,v})_{1\leq u,v\leq n_{1}}$
where $T_{h,u,v}:\operatorname{\mathbb{K}}^{m_{1}}\rightarrow\operatorname{\mathbb{K}}^{m_{1}}$ is linear. Since $\mathsf{H}_{1}$ and $\mathsf{H}_{2}$ commute, we have
$T_{h,u,v}\tau_{1}(g)=\tau_{1}(g)T_{h,u,v}$
for all $g\in\mathsf{H}_{1}$ and $1\leq u,v\leq n_{1}$. Thus by Schur’s lemma, $T_{h,u,v}=\lambda_{h,u,v}{\rm Id}_{\operatorname{\mathbb{K}}^{m_{1}}}$ for some $\lambda_{h,u,v}\in\operatorname{\mathbb{K}}$. Define a representation $\rho_{2}:\mathsf{H}_{2}\rightarrow\operatorname{\mathsf{SL}}(n_{1},\operatorname{\mathbb{K}})$ by $\rho_{2}(h)=(\lambda_{h,u,v})_{1\leq u,v\leq n_{1}}$.
Next we identify
$(\operatorname{\mathbb{K}}^{m_{1}})^{n_{1}}=\operatorname{\mathbb{K}}^{m_{1}}\otimes\operatorname{\mathbb{K}}^{n_{1}}$
by $(v_{1},\dots,v_{n_{1}})\mapsto v_{1}\otimes e_{1}+\cdots+v_{n_{1}}\otimes e_{n_{1}}$. Under this identification, $\mathsf{H}_{1}$ acts by
$h_{1}(v\otimes w)=(\tau_{1}(h_{1})(v))\otimes w$
and $\mathsf{H}_{2}$ acts by
$h_{2}(v\otimes w)=v\otimes(\rho_{2}(h_{2})(w)).$
Notice that if $V\subset\operatorname{\mathbb{K}}^{n_{1}}$ is a $\rho_{2}$-invariant linear subspace, then $\operatorname{\mathbb{K}}^{m_{1}}\otimes V\subset\operatorname{\mathbb{K}}^{m_{1}}\otimes\operatorname{\mathbb{K}}^{m_{2}}$ is $\mathsf{H}$-invariant. Hence $\rho_{2}$ is irreducible. Finally, setting $d_{1}=m_{1}$, $d_{2}=n_{1}$, and $\rho_{1}=\tau_{1}$ establishes the lemma. ∎
Using the previous lemma, we provide a sufficient condition for the real Zariski closure of an irreducible $\mathsf{P}_{k}$-Anosov subgroup to have no compact simple factors.
Lemma 4.5.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and $\mathsf{G}$ is the real Zariski closure of $\Gamma$. If either
  • $k=1$ or
  • $k>1$ and $\Gamma$ is also $\mathsf{P}_{k-1}$-Anosov,
then $\operatorname{\mathsf{G}}$ has no compact simple factors.
Proof.
It suffices to prove the lemma for the identity component $\operatorname{\mathsf{G}}^{0}$ of $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, $\Gamma\cap\operatorname{\mathsf{G}}^{0}$ is irreducible and its real Zariski closure is $\operatorname{\mathsf{G}}^{0}$. By part (1) of Lemma 4.3, $\operatorname{\mathsf{G}}^{0}$ has trivial center, so we may decompose $\operatorname{\mathsf{G}}^{0}$ into a direct product of simple Lie groups. In particular, we can write
$\operatorname{\mathsf{G}}^{0}=\mathsf{H}\times\mathsf{L}$
where $\mathsf{H}$ is the product of the non-compact simple factors and $\mathsf{L}$ is the product of the compact simple factors. Let $\mathsf{H}^{\prime}$ and $\mathsf{L}^{\prime}$ respectively denote the preimages of $\mathsf{H}$ and $\mathsf{L}$ under the quotient map $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{K}})$.
Suppose that $\mathsf{L}$ is not the trivial group. By Lemma 4.4, there exist integers $d_{1}$ and $d_{2}$ larger than $1$, representations $\rho_{1}:\mathsf{H}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(d_{1},\operatorname{\mathbb{K}})$ and $\rho_{2}:\mathsf{L}^{\prime}\rightarrow\operatorname{\mathsf{SL}}(d_{2},\operatorname{\mathbb{K}})$ and an identification
$\operatorname{\mathbb{K}}^{d}=\operatorname{\mathbb{K}}^{d_{1}}\otimes\operatorname{\mathbb{K}}^{d_{2}}$
such that an element $g=h\ell\in\mathsf{H}^{\prime}\mathsf{L}^{\prime}$ acts by
$g(v\otimes w)=(\rho_{1}(h)(v))\otimes(\rho_{2}(\ell)(w)).$
By conjugating we can assume that $\rho_{2}(\mathsf{L}^{\prime})\subset\mathsf{PU}(d_{2})$. Then endowing $\operatorname{\mathbb{K}}^{d_{1}}\otimes\operatorname{\mathbb{K}}^{d_{2}}$ with the inner product defined by
$\left\langle v_{1}\otimes w_{1},v_{2}\otimes w_{2}\right\rangle=\left\langle v_{1},v_{2}\right\rangle\left\langle w_{1},w_{2}\right\rangle,$
the associated singular values of an element $h\ell\in\mathsf{H}^{\prime}\mathsf{L}^{\prime}$ are
$\mu_{1}(\rho_{1}(h)),\dots,\mu_{1}(\rho_{1}(h)),\mu_{2}(\rho_{1}(h)),\dots,\mu_{2}(\rho_{1}(h)),\dots,\mu_{d_{1}}(\rho_{1}(h)),\dots,\mu_{d_{1}}(\rho_{1}(h))$
with each singular value repeated $d_{2}$ times. Hence $\Gamma$ cannot be $\operatorname{\mathsf{P}}_{1}$-Anosov or $\operatorname{\mathsf{P}}_{k-1,k}$-Anosov for some $k>1$. So we have a contradiction. ∎
4.3. Ruling out multiple non-compact simple factors
Next, we prove that if $\Gamma$ satisfies the hypothesis Theorem 4.1, then the real Zariski closure of $\Gamma$ cannot have multiple non-compact simple factors. To do so, we will use the following result.
Theorem 4.6.
([9, Thm. 1.5]) Suppose $\mathsf{H}$ is the identity component of a semisimple real algebraic Lie group and that $\Gamma<\mathsf{H}$ is $\mathsf{P}_{\theta}$-Anosov and Zariski dense. If $\Lambda_{\theta}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold of $\mathcal{F}_{\theta}$, then $\mathsf{P}_{\theta}$ is a maximal parabolic subgroup.
To apply Theorem 4.6 we first need to establish a few lemmas. The first is an immediate consequence of a result of Labourie [25, Prop. 10.3].
Lemma 4.7.
If $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible subgroup and $V_{1},\dots,V_{m},W_{1},\dots,W_{m}\subset\operatorname{\mathbb{K}}^{d}$ are linear subspaces with
$\dim V_{j}+\dim W_{j}\leq d$
for all $j=1,\dots,m$, then there exists $\gamma\in\Gamma$ where
$V_{j}\cap\gamma W_{j}=\{0\}$
for all $j=1,\dots,m$.
Proof.
Let $\mathsf{G}\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ denote the real Zariski closure of $\Gamma$ and let $\operatorname{\mathsf{G}}^{0}$ denote the connected component of the identity with respect to the Zariski topology in $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, $\operatorname{\mathsf{G}}^{0}$ is irreducible. By [25, Prop. 10.3] (this is stated for $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$, but the same arguments works for $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$) the set
$\mathcal{O}_{j}:=\{g\in\operatorname{\mathsf{G}}^{0}:V_{j}\cap gW_{j}=\{0\}\}$
is non-empty and also, by definition, open in the real Zariski topology. So $\mathcal{O}:=\cap_{j=1}^{m}\mathcal{O}_{j}$ is also non-empty and open in the real Zariski topology. Since $\Gamma\cap\operatorname{\mathsf{G}}^{0}$ is Zariski dense in $\operatorname{\mathsf{G}}^{0}$, then there exists some $\gamma\in\Gamma\cap\mathcal{O}$. ∎
The second guarantees the compactness of the subgroup of $\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ that fixes the limit set of a strongly irreducible $\operatorname{\mathsf{P}}_{k}$-Anosov subgroup.
Lemma 4.8.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and
$\mathsf{S}:=\{g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}}):gx=x\text{ for all }x\in\Lambda_{k}(\Gamma)\}.$
Then
  • If $k=1$, then $\mathsf{S}$ is trivial.
  • If $k>1$, then $\mathsf{S}$ is compact.
Proof.
When $k=1$, $\mathsf{S}$ is trivial, since $\Lambda_{1}(\Gamma)$ contains a projective frame (see [7, Lemma 2.18]) and an element of $\mathsf{PSL}(d,\mathbb{K})$ is determined by its action on a projective frame.
Now assume $k>1$. Suppose for a contradiction that $\mathsf{S}$ is not compact. Then we can find an unbounded sequence $\{s_{n}\}\subset\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})$ with $\{[s_{n}]\}\subset\mathsf{S}$. Let $s_{n}=m_{n}a_{n}\ell_{n}$ be a singular value decomposition, i.e $m_{n},\ell_{n}\in\mathsf{SU}(d,\operatorname{\mathbb{K}})$ and $a_{n}=\exp(\kappa(s_{n}))$. Since $\mathsf{SU}(d,\operatorname{\mathbb{K}})$ is compact, by passing to a subsequence we can suppose that $m_{n}\rightarrow m$ and $\ell_{n}\rightarrow\ell$. We can also assume that
$\frac{\sigma_{i_{1}}(s_{n})\cdots\sigma_{i_{k}}(s_{n})}{\sigma_{1}(s_{n})\cdots\sigma_{k}(s_{n})}\rightarrow a_{i_{1},\dots,i_{k}}\in[0,1]$
for all $1\leq i_{1}<\cdots<i_{k}\leq d$.
For any $i\in\{1,\dots,d\}$, let $\eta_{i}:\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})\to\operatorname{\mathsf{SL}}(\wedge^{i}\operatorname{\mathbb{R}}^{d})$ denote the representation induced by the linear action of $\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{R}})$ on $\wedge^{k}\operatorname{\mathbb{R}}^{d}$. The standard basis $\{e_{1},\dots,e_{d}\}$ of $\operatorname{\mathbb{K}}^{d}$ induces a standard basis
$\{e_{i_{1}}\wedge\cdots\wedge e_{i_{k}}:1\leq i_{1}<\dots<i_{k}\leq d\}$
of $\wedge^{k}\operatorname{\mathbb{K}}^{d}$. Then let $D_{i_{1},\dots,i_{k}}\in{\rm End}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ be the linear map that fixes the vector $e_{i_{1}}\wedge\dots\wedge e_{i_{k}}$ and sends all other standard basis vectors to 0. Observe that
$\frac{\eta_{k}(s_{n})}{\sigma_{1}(s_{n})\cdots\sigma_{k}(s_{n})}\rightarrow T:=\eta_{k}(m)\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq d}a_{i_{1},\dots,i_{k}}D_{i_{1},\dots,i_{k}}\right)\eta_{k}(\ell)$
in ${\rm End}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$. Further, if $x\in\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})\smallsetminus\operatorname{\mathbb{P}}(\ker T)$, then
$T(x)=\lim_{n\rightarrow\infty}\eta_{k}(s_{n})x.$ (5)
Now we view $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ as a subset of $\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ using the Plücker embedding, i.e. we identify $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ with the image of the embedding $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\to\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})$ given by $x\mapsto[h_{1}\wedge\dots\wedge h_{k}]$, where $\{h_{1},\dots,h_{k}\}$ is some (any) basis of $x$. Since $\mathsf{S}$, and hence $\eta_{k}(\mathsf{S})$, fixes $\Lambda_{k}(\Gamma)$ pointwise, Equation (5) implies that
$x=T(x)\quad\text{for all }\quad x\in\Lambda_{k}(\Gamma)\smallsetminus\operatorname{\mathbb{P}}(\ker T).$ (6)
Also, notice that $a_{1,\dots,k}=1$ and so
$\ker T$ $\subset\eta_{k}(\ell)^{-1}\{v\in\wedge^{k}\operatorname{\mathbb{R}}^{d}:v\wedge e_{k+1}\wedge\cdots\wedge e_{d}=0\}$ (7)
$=\{v\in\wedge^{k}\operatorname{\mathbb{R}}^{d}:v\wedge\eta_{d-k}(\ell)^{-1}(e_{k+1}\wedge\cdots\wedge e_{d})=0\}.$
Fix $x_{0}\in\Lambda_{k}(\Gamma)$, and choose vectors $v_{1},\dots,v_{k}\in\operatorname{\mathbb{K}}^{d}$ such that $x_{0}=[v_{1}\wedge\dots\wedge v_{k}]$. Since $\{s_{n}\}$ is unbounded, there exists $1\leq i_{1}^{\prime}<\cdots<i_{k}^{\prime}\leq d$ with $a_{i_{1}^{\prime},\dots,i_{k}^{\prime}}=0$. Then let $1\leq j_{1}<\cdots<j_{d-k}\leq d$ be the indices with
$\{i_{1}^{\prime},\dots,i_{k}^{\prime},j_{1},\dots,j_{d-k}\}=\{1,\dots,d\}.$
By Lemma 4.7 (applied with $W_{1}=W_{2}=\operatorname{Span}(v_{1},\dots,v_{k})$, $V_{1}=\ell^{-1}\operatorname{Span}(e_{k+1},\dots,e_{d})$, and $V_{2}=m\operatorname{Span}(e_{j_{1}},\dots,e_{j_{d-k}})$), there exists $\gamma\in\Gamma$ such that
$\eta_{k}(\gamma)(v_{1}\wedge\cdots\wedge v_{k})\wedge\eta_{d-k}(\ell)^{-1}(e_{k+1}\wedge\cdots\wedge e_{d})\neq 0$ (8)
and
$\eta_{k}(\gamma)(v_{1}\wedge\cdots\wedge v_{k})\wedge\eta_{d-k}(m)(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})\neq 0.$ (9)
Let $v:=\eta_{k}(\gamma)(v_{1}\wedge\dots\wedge v_{k})$ and let $x:=[v]=\eta_{k}(\gamma)(x_{0})$. By (7) and (8), $x\in\Lambda_{k}(\Gamma)\smallsetminus\operatorname{\mathbb{P}}(\ker T)$. Then by Equation (6), we have that $x=T(x)$, or equivalently, that $T(v)=\lambda v$ for some $\lambda\neq 0$. So
$v$ $\wedge\eta_{d-k}(m)(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})=\frac{1}{\lambda}T(v)\wedge\eta_{d-k}(m)(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})$
$=\frac{1}{\lambda}\det(m)\left(\left(\sum_{1\leq i_{1}<\cdots<i_{k}\leq d}a_{i_{1},\dots,i_{k}}D_{i_{1},\dots,i_{k}}\eta_{k}(\ell)(v)\right)\wedge(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})\right)$
$=\frac{1}{\lambda}\det(m)\left(a_{i_{1}^{\prime},\dots,i_{k}^{\prime}}D_{i_{1}^{\prime},\dots,i_{k}^{\prime}}\eta_{k}(\ell)(v)\right)\wedge(e_{j_{1}}\wedge\cdots\wedge e_{j_{d-k}})=0.$
This contradicts Equation (9). ∎
Lemma 4.9.
Suppose $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is a strongly irreducible $\mathsf{P}_{k}$-Anosov subgroup and $\mathsf{G}$ is the real Zariski closure of $\Gamma$. If $\Lambda_{k}(\Gamma)$ is a ${\mathcal{C}}^{1}$-submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$, then $\operatorname{\mathsf{G}}$ has exactly one non-compact simple factor.
Proof.
Let $\operatorname{\mathsf{G}}^{0}$ denote the identity component of $\operatorname{\mathsf{G}}$, and decompose
$\operatorname{\mathsf{G}}^{0}=\mathsf{H}\times\mathsf{L}$
where $\mathsf{H}$ is the product of the non-compact simple factors and $\mathsf{L}$ is the product of the compact simple factors. It suffices to prove that $\mathsf{H}$ is simple.
Let $\Gamma_{0}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$. By part (2) of Lemma 4.3, there is a non-empty subset $\theta\subset\Delta_{\operatorname{\mathsf{G}}^{0}}$ such that $\Gamma_{0}$ is $\operatorname{\mathsf{P}}_{\theta}$-Anosov, and a smooth embedding $\iota:\operatorname{\mathsf{G}}^{0}/\operatorname{\mathsf{P}}_{\theta}\to\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$ such that $\iota(\Lambda_{\theta}(\Gamma_{0}))=\Lambda_{k}(\Gamma)$. Since $\Lambda_{k}(\Gamma)$ is a smooth submanifold of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})$, $\Lambda_{\theta}(\Gamma_{0})$ is a smooth submanifold of $\mathsf{G}^{0}/\operatorname{\mathsf{P}}_{\theta}$, so Theorem 4.6 implies that $\mathsf{P}_{\theta}$ is a maximal parabolic subgroup.
Suppose for a contradiction that $\mathsf{H}=\prod_{j=1}^{m}\mathsf{H}_{j}$ has more than one simple factor. By the maximality of $\operatorname{\mathsf{P}}_{\theta}$, after relabeling the simple factors of $\mathsf{H}$, we can assume that
$\mathsf{P}_{\theta}\cap\mathsf{H}=\mathsf{Q}\times\mathsf{H}_{2}\times\cdots\times\mathsf{H}_{m}\times\mathsf{L},$
where $\mathsf{Q}$ is a parabolic subgroup of $\mathsf{H}_{1}$. It follows that the subgroup $\mathsf{H}_{2}\times\cdots\times\mathsf{H}_{m}\times L$ acts trivially $\mathsf{G}^{0}/\operatorname{\mathsf{P}}_{\theta}$, and hence also acts trivially on $\iota(\Lambda_{\theta}(\Gamma_{0}))=\Lambda_{k}(\Gamma)$. So by Lemma 4.8, $\mathsf{H}_{2}\times\cdots\times\mathsf{H}_{m}\times\mathsf{L}$ is compact which is a contradiction since each $\mathsf{H}_{j}$ is a non-compact simple factor. ∎
5. Rigidity
In this section, we complete the proof of Theorem 1.2 by establishing our rigidity result in the lattice case.
Theorem 5.1.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is a strongly irreducible (dk)(d-k)-hyperconvex subgroup and
$\delta^{\alpha_{k}}(\Gamma)=2,$
then $\Gamma$ is a uniform lattice in the image of an irreducible representation $\mathsf{PSL}(2,\operatorname{\mathbb{C}})\rightarrow\mathsf{PSL}(d,\operatorname{\mathbb{C}})$.
We originally had a more complicated proof of Theorem 5.1 that was more dynamical in flavor (a sketch of this argument is provided at the end of the section). We thank Andres Sambarino for suggesting the current simpler proof that uses Chow’s theorem [14].
Recall that a submanifold in a complex manifold is a complex submanifold if it is locally the graph of a holomorphic function. The following lemma is a main ingredient that we will need to apply Chow’s theorem.
Lemma 5.2.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (dk)(d-k)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{k}(\Gamma)$ is a complex submanifold of $\mathrm{Gr}_{k}(\mathbb{C}^{d})$.
Assuming Lemma 5.2, we will now prove Theorem 5.1.
Proof of Theorem 5.1.
Let $\mathsf{G}$ denote the real Zariski closure of $\Gamma$ and let $\operatorname{\mathsf{G}}^{0}$ denote the connected component of the identity in $\operatorname{\mathsf{G}}$. Since $\Gamma$ is strongly irreducible, it follows that $\Gamma^{\prime}:=\Gamma\cap\operatorname{\mathsf{G}}^{0}$, and hence also $\operatorname{\mathsf{G}}^{0}$, is irreducible.
Since $\delta^{\alpha_{k}}(\Gamma)=2$, by Corollary 3.2, $\Gamma$ is virtually isomorphic to a uniform lattice in $\mathsf{PSL}(2,\mathbb{C})$, so $\partial\Gamma^{\prime}=\partial\Gamma$ is a two-sphere. We may thus apply Lemma 5.2 to deduce that $\Lambda_{k}(\Gamma)$ is a complex submanifold in $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$, which is necessarily bilomorphic to $\mathbb{CP}^{1}$. We thus have an isomorphism of biholomorphism groups
$\mathsf{Aut}_{hol}(\Lambda_{k}(\Gamma))\cong\mathsf{Aut}_{hol}(\mathbb{CP}^{1})=\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
By Chow’s theorem [14], $\Lambda_{k}(\Gamma)$ is a complex (and also real) algebraic subvariety of $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$. Since $\operatorname{\mathsf{G}}$ is the real Zariski closure of $\Gamma$ and $\Lambda_{k}(\Gamma)$ is $\Gamma$-invariant, it follows that $\operatorname{\mathsf{G}}$ preserves $\Lambda_{k}(\Gamma)$. Since $\mathsf{G}\subset\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ acts on $\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{C}}^{d})$, and hence on $\Lambda_{k}(\Gamma)$, by biholomorphisms, we have a homomorphism
$q_{k}:\operatorname{\mathsf{G}}\rightarrow\mathsf{Aut}_{hol}(\Lambda_{k}(\Gamma))\cong\mathsf{PSL}(2,\operatorname{\mathbb{C}}).$
By Theorem 4.1, $\operatorname{\mathsf{G}}^{0}$ has no non-trivial proper normal subgroups, which in particular means that $q_{k}|_{\operatorname{\mathsf{G}}^{0}}$ is an isomorphism onto its image. Thus, the Lie algebra $\mathfrak{g}$ of $\mathsf{G}^{0}$ is isomorphic to a sub-algebra of $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$. Also, Theorem 4.1 implies that the Lie algebra $\mathfrak{g}$ is simple. Since $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$ has (real) rank $1$, and has six (real) dimensions, we deduce that either $\mathfrak{g}$ is isomorphic to $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$, or is simple, non-compact, rank $1$, and has dimension at most $5$. By the classification of rank $1$ simple Lie algebras, one deduces that $\mathfrak{g}$ is either isomorphic to $\mathfrak{sl}(2,\operatorname{\mathbb{C}})$ or to $\mathfrak{sl}(2,\operatorname{\mathbb{R}})$.
Since $\operatorname{\mathsf{G}}^{0}$ is connected and isomorphic to a subgroup of $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$, it now follows that $\operatorname{\mathsf{G}}^{0}$ is either isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ or $\mathsf{PSL}(2,\operatorname{\mathbb{R}})$. If $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{R}})$, then by the Milnor-Svarc lemma, $\partial\Gamma^{\prime}$ is a subset of $\mathbb{S}^{1}$, which contradicts the fact that $\partial\Gamma^{\prime}=\partial\Gamma$ is a $2$-sphere. Hence, we must have that $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
[…]
$\tau(g)[v_{1}\odot\dots\odot v_{d-1}]=[g(v_{1})\odot\dots\odot g(v_{d-1})]$
for all $v_{1},\dots,v_{d-1}\in\operatorname{\mathbb{C}}^{2}$ and all $g\in\mathsf{PSL}(2,\operatorname{\mathbb{C}})$. By Lemma 4.3, $\Gamma^{\prime\prime}:=\tau^{-1}(\Gamma^{\prime})$ is a $\operatorname{\mathsf{P}}_{\alpha_{1}}$-Anosov (i.e. convex cocompact) subgroup of $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ whose limit set if all of $\mathbb{CP}^{1}$, and hence is a lattice in $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$. It is straightforward to verify that for all $g\in\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ and $i\in\{1,\dots,d-1\}$,
$\alpha_{i}(\kappa(\tau(g)))=\alpha_{1}(\kappa(g)),$
where $\kappa$ on the left and the right are the Cartan projections for $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$ and $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$ respectively. It now follows from Theorem 2.2 that $\Gamma^{\prime}$, and hence $\Gamma$, is a $\operatorname{\mathsf{P}}_{\Delta}$-Anosov subgroup of $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$.
The map
$v]\in\mathbb{CP}^{1}\mapsto[v\odot\cdots\odot v]\in\operatorname{\mathbb{P}}({\rm Sym}^{d-1}(\operatorname{\mathbb{C}}^{2}))$
is a $\tau$-equivariant embedding that is a biholomorphism onto its image, which is the limit set $\Lambda_{1}(\Gamma)=\Lambda_{1}(\Gamma^{\prime})$. Further, $\operatorname{\mathsf{G}}$ acts by biholomorphisms on $\Lambda_{1}(\Gamma)$, so we have a homomorphism
$q_{1}:\operatorname{\mathsf{G}}\rightarrow\mathsf{Aut}_{hol}(\Lambda_{1}(\Gamma))$
from $\operatorname{\mathsf{G}}$ to the biholomorphism group of $\Lambda_{1}(\Gamma)$. Since $\Lambda_{1}(\Gamma)$ contains $(d+1)$-points in general position, the only element of $\operatorname{\mathsf{G}}$ acting trivially on $\Lambda_{1}(\Gamma)$ is the identity. Hence $\ker q_{1}$ is trivial. At the same time, since $\operatorname{\mathsf{G}}^{0}$ is isomorphic to $\mathsf{PSL}(2,\operatorname{\mathbb{C}})\cong\mathsf{Aut}_{hol}(\Lambda_{1}(\Gamma))$, $q_{1}|_{\operatorname{\mathsf{G}}^{0}}$ is an isomorphism. Thus $\operatorname{\mathsf{G}}=\operatorname{\mathsf{G}}^{0}$. ∎
It now remains to prove Lemma 5.2. To do so, we first observe from the following result of Pozzetti–Sambarino–Wienhard [30] (see also Zhang–Zimmer [37] for the case when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{R}}$) that $\Lambda_{k}(\Gamma)$ is a differentiable submanifold.
Theorem 5.3.
([30, Prop. 7.1]) Suppose that $\Gamma\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is (d1)(d-1)-hyperconvex subgroup. If $\{(z_{n},w_{n})\}$ is a sequence of distinct points in $\Lambda_{\theta_{1}}(\Gamma)$ that converges to $(x,x)$, then
$z_{n}^{1}\oplus y_{n}^{1}\to x^{2}.$
Notice that for every $x\in\Lambda_{\theta_{1}}(\Gamma)$, the set of points in $\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ that lie in $x^{2}$ is a projective subspace of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ that we also denote by $x^{2}$. Furthermore, for any affine chart $U\subset\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ that contains $x^{1}$, $x^{2}\cap U$ is an affine subspace of $U$ that contains $x^{1}$, so we may view $x^{2}$ as a $\operatorname{\mathbb{K}}$-linear subspace of $T_{x^{1}}\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})\cong T_{x^{1}}U$. Thus, in the case that $\operatorname{\mathbb{K}}=\mathbb{C}$ and $\partial\Gamma$ is a two-sphere, Theorem 5.3 immediately implies the following consequence. (For comparison, see [30, Prop. 7.4] and [37, Thm. 1.1].)
Corollary 5.4.
If $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (d1)(d-1)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{1}(\Gamma)$ is a $\mathcal{C}^{1}$-submanifold of $\mathbb{P}(\mathbb{C}^{d})$. Furthermore, for every $x^{1}\in\Lambda_{1}(\Gamma)$, and the tangent space $T_{x^{1}}\Lambda_{1}(\Gamma)\subset T_{x^{1}}\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{d})$ is a complex linear subspace.
To further upgrade $\Lambda_{1}(\Gamma)$ to a complex submanifold of $\operatorname{\mathbb{P}}(\operatorname{\mathbb{C}}^{d})$, we use the following general statement.
Lemma 5.5.
Suppose $N$ is a ${\mathcal{C}}^{1}$-submanifold of a complex manifold $M$. If for every $x\in N$ the tangent space $T_{x}N$ is a complex linear subspace of $T_{x}M$, then $N$ is a complex submanifold. In particular, if $\Gamma\subset\mathsf{PSL}(d,\mathbb{C})$ is (d1)(d-1)-hyperconvex subgroup and $\partial\Gamma$ is a two-sphere, then $\Lambda_{1}(\Gamma)$ is a complex submanifold of $\mathbb{P}(\mathbb{C}^{d})$.
Proof.
Since being a complex manifold is a local condition, it suffices to prove the following: If $U\subset\operatorname{\mathbb{C}}^{n}$ is an open connected set and $f:U\rightarrow\operatorname{\mathbb{C}}^{m}$ is a ${\mathcal{C}}^{1}$-smooth function such that every tangent plane to the ${\mathcal{C}}^{1}$-submanifold
$N:=\{(z,f(z))\in\operatorname{\mathbb{C}}^{m+n}:z\in U\}$
is complex linear, then $f$ is holomorphic. Now if $x=(z,f(z))\in N$, then
$T_{x}N=\{(v,D(f)_{z}v)\in T_{x}\operatorname{\mathbb{C}}^{m+n}:v\in T_{z}\operatorname{\mathbb{C}}^{n}\}.$
Further, by hypothesis,
$(iv,iD(f)_{z}v)=i\cdot(v,D(f)_{z}v)\in T_{x}N$
for all $v\in T_{z}\operatorname{\mathbb{C}}^{n}$. So we must have $iD(f)_{z}v=D(f)_{z}iv$ for all $z\in U$ and $v\in T_{z}\operatorname{\mathbb{C}}^{n}$, i.e. $f$ is holomorphic. ∎
We may now assemble the above results to prove Lemma 5.2.
Proof of Lemma 5.2.
By Proposition 2.6, it suffices to prove the lemma for the case when $k=1$. The case when $k=1$ is immediate from Corollary 5.4 and Lemma 5.5. ∎
Sketch of a more dynamical proof
We will now give a sketch of our original, more dynamical proof prior to the simplification suggested by Andres Sambarino.
For any $z\in\Lambda_{\theta_{k}}(\Gamma)$, Farre–Pozzetti–Viaggi [19] defined a homeomorphism
$\Lambda_{k}(\Gamma)\to\mathbb{P}\left(z^{d-k+1}/z^{d-k-1}\right)\cong\mathbb{CP}^{1}$
called the tangent projection at $z$, and verified that it is bilipschitz on all compact subsets of $\Lambda_{k}(\Gamma)-\{z^{k}\}$. Via this homeomorphism, the $\Gamma$-action on $\Lambda_{k}(\Gamma)$ induces a $\Gamma$-action
$\rho_{z}:\Gamma\to{\rm Homeo}(\mathbb{CP}^{1})$
on $\mathbb{CP}^{1}$, which they then prove is uniformly quasiconformal. A theorem of Sullivan [33] then implies that there exists a a uniform lattice $\hat{\Gamma}\subset\mathsf{PSL}(2,\mathbb{C})$ and a quasi-conformal conjugacy $\phi:\rho_{z}(\Gamma)\to\hat{\Gamma}$.
By Theorem 3.1, the $\alpha_{k}$-Patterson–Sullivan measure for $\Gamma$ on $\Lambda_{k}(\Gamma)$ is Ahlfors regular and the $\alpha_{1}$-Patterson–Sullivan measure of $\hat{\Gamma}$ on $\mathbb{CP}^{1}$ is mutually absolutely continuous with the Lebesgue measure. Since quasi-conformal maps preserve the Lebesgue measure class, we can then conclude that their pull-backs to $\partial\Gamma$ are mutually absolutely continuous. Then a result of Blayac–Canary–Zhu–Zimmer [5, Thm. 13.2] (see also Sambarino [32, Cor. 3.3.4]) implies that
$\alpha_{k}(\nu(\gamma))=\alpha_{1}(\nu(\phi(\rho_{z}(\gamma))))$ (10)
for all $\gamma\in\Gamma$, where $\nu$ on the left denotes the Jordan projection in $\mathsf{PSL}(d,\operatorname{\mathbb{C}})$, while $\nu$ on the right denotes the Jordan projection for $\mathsf{PSL}(2,\operatorname{\mathbb{C}})$.
We then consider the identity component $\mathsf{G}^{0}$ of the real Zariski closure $\mathsf{G}$ of $\Gamma$. We know, by Theorem 4.1, that $\mathsf{G}^{0}$ is simple. We may now use an argument of Dal’bo–Kim [15] to show that $\mathsf{G}^{0}$ is isomorphic to $\mathsf{PSL}(2,\mathbb{C})$. Once we have done so, we can complete the proof of Theorem 5.1 just as above.
Finally, we briefly sketch the argument that $\mathsf{G}^{0}$ is isomorphic to $\mathsf{PSL}(2,\mathbb{C})$. Consider the subgroup
$H_{0}:=\{(\gamma,\phi(\rho_{x}(\gamma)):\gamma\in\Gamma\cap\operatorname{\mathsf{G}}^{0}\}\subset\mathsf{G}^{0}\times\mathsf{PSL}(2,\mathbb{C}),$
and let $Z$ be its real Zariski closure in $\mathsf{G}^{0}\times\mathsf{PSL}(2,\mathbb{C}).$ Equation (10) and a result of Benoist [3] implies that $H_{0}$ is not Zariski dense, so $Z\neq\mathsf{G}^{0}\times\mathsf{PSL}(2,\mathbb{C}).$ Consider the projection maps $\pi_{1}:Z\to\mathsf{G}^{0}$ and $\pi_{2}:Z\to\mathsf{PSL}(2,\mathbb{C})$, each of which is surjective. The kernel of $\pi_{1}$ is normal in $\{1\}\times\mathsf{PSL}(2,\mathbb{C})$ and cannot be all of $\{1\}\times\mathsf{PSL}(2,\mathbb{C})$. Since $\mathsf{PSL}(2,\mathbb{C})$ is simple and has trivial center, $\pi_{1}$ is an isomorphism. Similarly, since $\mathsf{G}^{0}$ is simple and has trivial center, $\pi_{2}$ is an isomorphism. Therefore, $\pi_{1}\circ\pi_{2}^{-1}$ is an isomorphism from $\mathsf{G}^{0}$ is isomorphic to $\mathsf{PSL}(2,\mathbb{C})$.
Appendix A Reducing to the projective case
In this appendix we give the proof of Proposition 2.6.
For the rest of the section fix $k\in\{1,\dots,d-1\}$ and, as usual, let $(e_{1},\dots,e_{d})$ denote the standard basis on $\operatorname{\mathbb{K}}^{d}$. We identify
$\operatorname{\mathbb{K}}^{\bar{d}}=\wedge^{k}\operatorname{\mathbb{K}}^{d},\quad\bar{d}:={d\choose k}=\dim\wedge^{k}\operatorname{\mathbb{K}}^{d},$
using the basis
$(e_{i_{1}}\wedge\dots\wedge e_{i_{k}})_{1\leq i_{1}<\dots<i_{k}\leq d}$
where the order is in lexicographic order of the subscripts. This gives an identification $\mathsf{PSL}(\wedge^{k}\operatorname{\mathbb{K}}^{d})=\mathsf{PSL}(\bar{d},\operatorname{\mathbb{K}})$. There is a natural representation $\tilde{\eta}:\operatorname{\mathsf{SL}}(d,\operatorname{\mathbb{K}})\to\operatorname{\mathsf{SL}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})=\operatorname{\mathsf{SL}}(\bar{d},\operatorname{\mathbb{K}})$ defined by
$\tilde{\eta}(g)(v_{1}\wedge\dots\wedge v_{k})=g(v_{1})\wedge\dots\wedge g(v_{k}),$
which descends to a representation
$\eta:\mathsf{PSL}(d,\operatorname{\mathbb{K}})\to\mathsf{PSL}(\wedge^{k}\operatorname{\mathbb{K}}^{d})=\mathsf{PSL}(\bar{d},\operatorname{\mathbb{K}}).$
Observe then that the smooth embedding $E_{k}:\operatorname{\mathrm{Gr}}_{k}(\operatorname{\mathbb{K}}^{d})\hookrightarrow\operatorname{\mathbb{P}}(\wedge^{k}\operatorname{\mathbb{K}}^{d})=\operatorname{\mathbb{P}}(\operatorname{\mathbb{K}}^{\bar{d}})$ defined by
$E_{k}(\operatorname{Span}(v_{1},\dots,v_{k}))=[v_{1}\wedge\cdots\wedge v_{k}]$
is $\eta_{k}$-equivariant, and is holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$.
Set $\bar{\Delta}:=\{1,\dots,\bar{d}-1\}$ and $\bar{\theta}_{1}:=\{1,2,\bar{d}-2,\bar{d}-1\}\subset\bar{\Delta}$. For any non-empty subset $\theta\subset\Delta$ (respectively, $\bar{\theta}\subset\bar{\Delta}$), let $\pi_{\theta}:\operatorname{\mathcal{F}}_{\Delta}(\operatorname{\mathbb{K}}^{d})\to\operatorname{\mathcal{F}}_{\theta}(\operatorname{\mathbb{K}}^{d})$ (respectively, $\bar{\pi}_{\bar{\theta}}:\operatorname{\mathcal{F}}_{\bar{\Delta}}(\operatorname{\mathbb{K}}^{\bar{d}})\to\operatorname{\mathcal{F}}_{\bar{\theta}}(\operatorname{\mathbb{K}}^{\bar{d}})$) denote the forgetful projection. We record the following well-known properties of $\eta$ that we will use.
  1. (I)
    The representation $\eta$ sends $\operatorname{\mathsf{P}}_{\Delta}\subset\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ to $\operatorname{\mathsf{P}}_{\bar{\Delta}}\subset\mathsf{PSL}(\bar{d},\operatorname{\mathbb{K}})$, so we may define an $\eta$-equivariant, smooth (holomorphic when $\operatorname{\mathbb{K}}=\operatorname{\mathbb{C}}$) embedding
    $E:\operatorname{\mathcal{F}}_{\Delta}(\operatorname{\mathbb{K}}^{d})\to\operatorname{\mathcal{F}}_{\bar{\Delta}}(\operatorname{\mathbb{K}}^{\bar{d}})\quad\text{given by}\quad E(g\operatorname{\mathsf{P}}_{\Delta})=\eta(g)\operatorname{\mathsf{P}}_{\bar{\Delta}}.$
  2. (II)
    For any $x\in\operatorname{\mathcal{F}}_{\Delta}(\operatorname{\mathbb{K}}^{d})$, we have
    $E(x)^{1}$ $=\operatorname{Span}(\{v_{1}\wedge\dots\wedge v_{k}:\operatorname{Span}(v_{1},\dots,v_{k})=x^{k}\}),$
    $E(x)^{2}$ $=\operatorname{Span}(\{v_{1}\wedge\dots\wedge v_{k}:x^{k-1}\subset\operatorname{Span}(v_{1},\dots,v_{k})\subset x^{k+1}\}),$
    $E(x)^{\bar{d}-2}$ $=\operatorname{Span}(\{v_{1}\wedge\dots\wedge v_{k}:\operatorname{Span}(v_{1},\dots,v_{k})\cap x^{d-k-1}\neq 0\}$
    $\hskip 28.45274pt\cup\{v_{1}\wedge\dots\wedge v_{k}:\dim(\operatorname{Span}(v_{1},\dots,v_{k})\cap x^{d-k+1})\geq 2\}),$
    $E(x)^{\bar{d}-1}$ $=\operatorname{Span}(\{v_{1}\wedge\dots\wedge v_{k}:\operatorname{Span}(v_{1},\dots,v_{k})\cap x^{d-k}\neq 0\}).$
    In particular, $E(x)^{1}=E_{k}(x^{k})$ and we have well-defined, $\tau$-equivariant, smooth embeddings
    $E_{\theta_{k}}:\operatorname{\mathcal{F}}_{\theta_{k}}(\operatorname{\mathbb{K}}^{d})\to\operatorname{\mathcal{F}}_{\bar{\theta}_{1}}(\operatorname{\mathbb{K}}^{\bar{d}}).$
    for any $x\in\operatorname{\mathcal{F}}_{\Delta}(\operatorname{\mathbb{K}}^{d})$.
  3. (III)
    The embedding $E$ preserves transversality of pairs of flags, so the same is true for $E_{\theta_{k}}$.
  4. (IV)
    The representation $\tilde{\eta}$ sends $\mathsf{SU}(d,\operatorname{\mathbb{K}})$ into $\mathsf{SU}(\bar{d},\operatorname{\mathbb{K}})$.
  5. (V)
    If $a={\rm Diag}(a_{1},\dots,a_{d})\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ is real valued, then $\tilde{\eta}(a)$ is a diagonal, real valued matrix whose diagonal entries are all possible products of $k$ of the $d$ diagonal entries of $a$, and whose $(1,1)$ and $(2,2)$ entries are
    $\prod_{i=1}^{k}a_{i}\quad\text{and}\quad a_{k+1}\prod_{i=1}^{k-1}a_{i}$
    respectively. Furthermore, if $a_{1}\geq\dots\geq a_{d}$, then the largest and second largest entries of $\tilde{\eta}(a)$ are respectively its $(1,1)$ and $(2,2)$ entries, and its third largest entry is
    $\left\{\begin{array}[]{ll}\displaystyle a_{k+1}a_{k}\prod_{i=1}^{k-2}a_{i}&\text{if }d=k+1,\\ \displaystyle\max\left\{a_{k+2}\prod_{i=1}^{k-1}a_{i},\,a_{k+1}a_{k}\prod_{i=1}^{k-2}a_{i}\right\}&\text{if }d\geq k+2.\end{array}\right$.
  6. (VI)
    By (I), (IV), and (V), if $g=man$ is the Iwasawa decomposition of $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$, then $\eta(g)=\eta(m)\eta(a)\eta(n)$ is the Iwasawa decomposition of $\eta(g)$, and $\alpha_{1}(\log\tau(a))=\alpha_{k}(\log a)$. In particular, if $\bar{B}:\mathsf{PSL}(\bar{d},\operatorname{\mathbb{K}})\times\operatorname{\mathcal{F}}_{\bar{\Delta}}\to\operatorname{\mathbb{R}}$ is the Iwasawa cocycle of $\mathsf{PSL}(\bar{d},\operatorname{\mathbb{K}})$, then for all $g\in\mathsf{PSL}(d,\operatorname{\mathbb{K}})$ and $x\in\operatorname{\mathcal{F}}_{\Delta}(\operatorname{\mathbb{K}}^{d})$, we have
    $\alpha_{1}(\bar{B}(\eta(g),E(x))=\alpha_{k}(B(g,x)).$
  7. (VII)
    By (IV) and (V), if $g=m_{1}e^{\kappa(g)}m_{2}$ is a singular value decomposition of $g$, then there is a permutation matrix $p\in\mathsf{PGL}(\bar{d},\operatorname{\mathbb{K}})$ whose $(1,1)$, $(2,2)$, $(\bar{d}-1,\bar{d}-1)$, and $(\bar{d},\bar{d})$ entries are all $1$ such that $\eta(g)=\bar{m}_{1}e^{\kappa(\eta(g))}\bar{m}_{2}$ is a singular value decomposition of $\eta(g)$, where $\bar{m}_{1}:=\eta(m_{1})p^{-1}$ and $\bar{m}_{2}:=p\eta(m_{2})$. From this, we may deduce that $\alpha_{1}(\kappa(\eta(g)))=\alpha_{k}(\kappa(g))$,
    $\alpha_{2}(\kappa(\eta(g)))=\left\{\begin{array}[]{ll}\alpha_{k-1}(\kappa(g))&\text{if }d=k+1,\\ \max\{\alpha_{k-1}(\kappa(g)),\alpha_{k+1}(\kappa(g))\}&\text{if }d\geq k+2,\end{array}\right$.
    and $E_{\theta_{k}}(\operatorname{\mathsf{U}}_{\theta_{k}}(g))=\operatorname{\mathsf{U}}_{\bar{\theta}_{1}}(\eta(g))$.
Proof of Proposition 2.6.
First, we prove (1). By (VII), the fact that $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-divergent implies that $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\bar{\theta}_{1}}$-divergent, and $\Lambda_{\bar{\theta}_{1}}(\eta(\Gamma))=E_{\theta_{k}}(\Lambda_{\theta_{k}}(\Gamma))$. In particular, $\Lambda_{1}(\eta(\Gamma))=E_{k}(\Lambda_{k}(\Gamma))$. Since $\Gamma$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-transverse, (III) implies that $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\bar{\theta}_{1}}$-transverse. Since $E_{\theta_{k}}$ is an $\eta$-equivariant embedding, the fact that every point in $\Lambda_{\theta_{k}}(\Gamma)$ is a conical point for the $\Gamma$-action implies that every point in $\Lambda_{\bar{\theta}_{1}}(\eta(\Gamma))$ is a conical point for the $\tau(\Gamma)$-action. Hence, $\eta(\Gamma)$ is $\operatorname{\mathsf{P}}_{\theta_{k}}$-Anosov.
Let $x,y,z\in\Lambda_{\theta_{k}}(\Gamma)$ be any mutually distinct triples of flags. Since $x$ and $z$ are transverse, we may choose a basis $(v_{1},\dots,v_{d})$ of $\operatorname{\mathbb{K}}^{d}$ such that
$x^{k-1}=\operatorname{Span}(v_{1},\dots,v_{k-1}),\quad x^{k}\cap z^{d-k+1}=\operatorname{Span}(v_{k}),$
$x^{k+1}\cap z^{d-k}=\operatorname{Span}(v_{k+1}),\quad\text{and}\quad z^{d-k-1}=\operatorname{Span}(v_{k+2},\dots,v_{d}).$
For any tuple ${\bf j}=(j_{1},\dots,j_{k})$ of integers such that $1\leq j_{1}<\dots<j_{k}\leq d$, set $v_{\bf j}:=v_{j_{1}}\wedge\dots\wedge v_{j_{k}}$. Then observe from the description of $E_{\theta_{k}}$ given in (II) that
$E_{\theta_{k}}(x)^{1}=\operatorname{Span}(v_{(1,\dots,k)})\quad\text{and}\quad E_{\theta_{k}}(z)^{\bar{d}-2}=\operatorname{Span}(\{v_{\bf j}:{\bf j}\neq(1,\dots,k),(1,\dots,k-1,k+1)\}).$
Also, since $\Gamma$ is (dk)(d-k)-hyperconvex, we have
$\Big((x^{k}\cap z^{d-k+1})+z^{d-k-1}\Big)\cap\Big((y^{k}\cap z^{d-k+1})+z^{d-k-1}\Big)=z^{d-k-1}.$
Since $v_{k}\in(x^{k}\cap z^{d-k+1})+z^{d-k-1}$ and $v_{k}\notin z^{d-k-1}$, we must have
$v_{k}\notin(y^{k}\cap z^{d-k+1})+z^{d-k-1}.$
Hence, there is some $a_{k},a_{k+2},\dots,a_{d}\in\operatorname{\mathbb{R}}$ such that
$y^{k}\cap z^{d-k+1}=\operatorname{Span}\left(v_{k+1}+a_{k}v_{k}+\sum_{j=k+2}^{d}a_{j}v_{j}\right).$
Since $y^{k-1}+z^{d-k+1}=\operatorname{\mathbb{K}}^{d}$, it now follows that
$E_{\theta_{k}}(y)^{1}=\operatorname{Span}\left(v_{(1,\dots,k-1,k+1)}+\sum_{{\bf j}\neq(1,\dots,k-1,k+1)}b_{\bf j}v_{\bf j}\right)$
for some $\{b_{\bf j}\}\subset\operatorname{\mathbb{K}}$. Thus, $E_{\theta_{k}}(x)^{1}+E_{\theta_{k}}(y)^{1}+E_{\theta_{k}}(z)^{\bar{d}-2}=\operatorname{\mathbb{K}}^{\bar{d}}$. Since $x,y,z$ was arbitrary, it follows that $\eta(\Gamma)$ is $(\bar{d}-1)$-hyperconvex. This finishes the proof of (1).
Statements (2) and (3) are immediate consequences of (VII) and (VI) respectively. ∎